### 18 Mar 2026 v2

# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes. We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}\qquad(x\to\infty),$
and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$. We have not found this full-support statement in the literature. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma and is moreover continuous and purely singular. We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1. As a simple conservative explicit bound, for any prime $q>2^{30}$ the expected overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [1], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. The fraction of such matrices is denoted $c_{n}(p)$; since it is decreasing in $n$, the dimension-uniform constant governing the reduction from decision-LWE to decision-prim-lwe is $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$. The precise formulas are recorded in Section 2. Sehrawat, Yeo, and Desmedt [1, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally; see Section 2 for the precise statement.
Main results
We resolve this question and establish several natural refinements.
1. Unconditional vanishing (Theorem 3). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
2. Sharp order of the minimum (Proposition 1, Corollary 6). By additionally invoking Linnik’s theorem on the least prime in an arithmetic progression, together with standard prime-number-theorem consequences for primorials, we establish
$\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 8. Hence the decay to zero is completely quantified and extremely slow. The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$, well within practical ranges. The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
3. Limiting distribution (Theorems 6 and 7, Corollary 2). The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular. If $G$ denotes the distribution function, then
$0<G(\tau)<1\qquad(0<\tau<1/2),$
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$. The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4]. Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p. 263]; we include a self-contained proof for the reader’s convenience. Our new contributions on the distributional side are twofold. First, we establish that the support of the limiting law is exactly $0,\tfrac{1}{2}]$ by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Second, we transfer the classical shifted-prime law for $\varphi(p-1)/(p-1)$ to the prim-lwe density $c(p)$ via a transport lemma (Lemma 5).
4. Explicit bounds for cryptographic moduli (Theorem 8, Table 1). We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure. The following consequence singles out the cryptographic implication.
Corollary 1 (Cryptographic consequence).
Let
$P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779.$
Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then, in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ satisfies
$\frac{1}{c(q)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
For each fixed dimension $n$, the exact expected overhead is $1/c_{n}(q)\leq 1/c(q)$.
The general $\omega(q-1)$-based bound is immediate from Theorem 8; the sharper NIST figures use the exact factorizations recorded in Section 6.
Related work
A related distribution problem was studied by Aivazidis and Sofos [6], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in prim-lwe. On the distributional side, Deshouillers and Hassani [4, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [4, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [4].
Organization
Section 2 records the precise definitions and the open question from [1]. Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$. Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$. Section 5 develops the limiting distribution theory. Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$. Section 7 discusses the implications for prim-lwe.
Notation
We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order. All limits are taken as $x\to\infty$ unless stated otherwise. We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively. $A\asymp B$ means both $A\ll B$ and $A\gg B$. We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm. We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e. an implied constant of $1$.
2 Background and the Open Question
We now record the precise setup from [1] and state the open question. Following [1], define
$M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
The determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{p})\longrightarrow\mathbb{F}_{p}^{\times}$
is a surjective homomorphism, so all its fibres have the same cardinality. Therefore a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
it follows that
$c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
recovering [1, Corollary 4]. Consequently, the limit
$c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
exists and is strictly positive for each fixed $p$. The quantity $c_{n}(p)$ is the exact acceptance probability in dimension $n$. Since $c_{n}(p)$ is decreasing in $n$, the dimension-uniform constant that enters the prim-lwe reduction is
$\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)=c(p).$
Thus $c(p)$ is not the exact finite-dimensional density but the limiting, dimension-uniform reduction constant.
Recall that one may equivalently bundle $n$ standard LWE secret vectors into a matrix and view decision-LWE in matrix form: given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
$(\mathbf{A},\,\mathbf{A}\mathbf{S}+\mathbf{E})$
from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7]. Via the identification $\mathbb{Z}_{q}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
In [1, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes
$p=\prod_{i=1}^{k}p_{i}+1,$
then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{k}\frac{p_{i}-1}{p_{i}}\longrightarrow 0$
by Mertens’ theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [8] for computational results on primorial primes.
3 Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results.
Theorem 1 (Dirichlet [9], 1837).
For coprime integers $a$ and $q$, there are infinitely many primes $p\equiv a\pmod{q}$.
Theorem 2 (Mertens’ Third Theorem [10], 1874).
As $x\to\infty$,
$\prod_{\begin{subarray}{c}q\leq x\\ q\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
We also record the following simple observation.
Lemma 1 (Lower bound for the matrix-count Euler product).
For all primes $p\geq 3$,
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;\frac{1}{2}.$
Proof.
For each $J\geq 1$, set $a_{j}:=p^{-j}$. Then $0<a_{j}\leq 1$ for all $j$, so the finite-product inequality
$\prod_{j=1}^{J}(1-a_{j})\geq 1-\sum_{j=1}^{J}a_{j}$
gives
$\prod_{j=1}^{J}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{J}\frac{1}{p^{j}}.$
Letting $J\to\infty$ yields
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;1-\sum_{j=1}^{\infty}\frac{1}{p^{j}}=1-\frac{1}{p-1}\geq\frac{1}{2}\qquad(p\geq 3).$
Theorem 3 (Unconditional vanishing of the infimal PRIM-LWE density).
$\inf_{p\textup{ prime}}c(p)=0,\qquad\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0.$
Proof.
For each $k\geq 1$ define the $k$th primorial
$N_{k}\;=\;\prod_{i=1}^{k}p_{i}.$
Since $\gcd(1,N_{k})=1$, Theorem 1 guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $q\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
where the inequality holds because each factor $\bigl(1-\frac{1}{q}\bigr)$ lies in $(0,1)$, so the full product over all prime divisors $q\mid(p-1)$ is no larger than the partial product over the subset $p_{1},\dots,p_{k}$.
By Mertens’ third theorem (Theorem 2),
$\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since the infinite product satisfies $\prod_{j=1}^{\infty}(1-p^{-j})\leq 1$ for every prime $p$, we obtain
$c(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since for every $k$ there are infinitely many primes $p\equiv 1\pmod{N_{k}}$, the displayed upper bound holds for arbitrarily large primes $p$. Letting $k\to\infty$ gives
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0,$
and the infimum statement follows immediately. ∎
4 Quantitative Refinements
We first record the form of Linnik’s theorem used in the proof.
Theorem 4 (Least prime in an arithmetic progression; Linnik [11]).
There exist absolute constants $C_{0},L>0$ such that for every modulus $q\geq 1$ and every residue class $a\pmod{q}$ with $\gcd(a,q)=1$, the least prime $p(a,q)\equiv a\pmod{q}$ satisfies
$p(a,q)\leq C_{0}q^{L}.$
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [12]).
Let
$N_{k}=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p.$
Then
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}\qquad(k\to\infty).$
Consequently, if
$m(x):=\max\{k\geq 1:N_{k}\leq x\},$
then
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x\qquad(x\to\infty).$
Proof.
The first two assertions are standard consequences of the prime number theorem. Since
$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1},$
we have
$\log N_{m(x)}\leq\log x<\log N_{m(x)}+\log p_{m(x)+1}.$
Now
$\log p_{m(x)+1}=o\!\bigl(p_{m(x)}\bigr)=o\!\bigl(\log N_{m(x)}\bigr),$
because $\log N_{m(x)}=\vartheta(p_{m(x)})\sim p_{m(x)}$ and $p_{m(x)+1}\sim p_{m(x)}$. Hence $\log x\sim\log N_{m(x)}$, and therefore $p_{m(x)}\sim\log x$. ∎
Proposition 1 (Upper bound for the minimum density up to $x$).
There exists an absolute constant $C>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;\frac{C}{\log\log x}.$
Proof.
By Theorem 4, fix absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $q$ is at most $C_{0}q^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that
$C_{0}N_{k}^{\,L}\leq x.$
Then
$p^{*}(k)\leq C_{0}N_{k}^{\,L}\leq x.$
Such a $k$ exists for all sufficiently large $x$, and $k\to\infty$ as $x\to\infty$.
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x,$
and hence $p_{k}\ll\log x$. Therefore $\log p_{k}=O(\log\log x)$.
Conversely, the maximality of $k$ forces $C_{0}N_{k+1}^{\,L}>x$, so
$L\log N_{k+1}>\log x-\log C_{0}.$
Since $\log N_{k+1}=\log N_{k}+\log p_{k+1}$, while
$\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}\quad\text{and}\quad\log p_{k+1}\sim\log p_{k}=o(p_{k})=o(\log N_{k}),$
we have $\log N_{k+1}\sim\log N_{k}$. Hence $\log N_{k}\gg\log x$, so $p_{k}\gg\log x$ and therefore $\log p_{k}=\Omega(\log\log x)$. Combined with the upper bound just proved, this yields
$\log p_{k}\asymp\log\log x.$
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 3 gives
$c\bigl(p^{*}(k)\bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\asymp\;\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;c\bigl(p^{*}(k)\bigr)\;\ll\;\frac{1}{\log\log x},$
as claimed. ∎
5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp. 257–258] and Deshouillers–Hassani [4].
Our contributions in this section are:
  1. (i)
    The full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
  2. (ii)
    The transport of the classical shifted-prime law to the prim-lwe density $c(p)$ via Lemma 5, establishing that $c(p)$ shares the same limiting distribution as $\varphi(p-1)/(p-1)$.
We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9.
5.1 The distribution of $\varphi(p-1)/(p-1)$
One can bypass the general Erdős–Wintner machinery and use the following direct p1p-1-specific input from Deshouillers and Hassani.
Lemma 3 (Erdős–Wintner hypotheses for the totient function).
Let $h$ be the strongly additive function $h(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<h(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad h(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
Consequently the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while $\sum_{\ell}h(\ell)/\ell$ and $\sum_{\ell}h(\ell)^{2}/\ell$ both converge. Since $h(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
The same conclusions hold for $g(n)=\varphi(n)/n$ via $\log g(\ell)=-h(\ell)$.
Proposition 2 (Deshouillers–Hassani; special case of [4, Proposition 2.1 and equation (2.1)]).
For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\text{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
Proof.
Apply Deshouillers and Hassani [4, Proposition 2.1] to
$g(n):=\frac{\varphi(n)}{n}.$
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [4, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
$\sum_{\begin{subarray}{c}\ell\equiv 1\pmod{m}\\ g(\ell)\neq 1\end{subarray}}\frac{1}{\ell}$
diverges. Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
$\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
by the divergence of the prime harmonic series in every reduced residue class. Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
If $p\equiv 1\pmod{m}$, then every prime divisor of $m$ also divides p1p-1, so
$\frac{\varphi(p-1)}{p-1}=\prod_{q\mid(p-1)}\Bigl(1-\frac{1}{q}\Bigr)\leq\prod_{q\mid m}\Bigl(1-\frac{1}{q}\Bigr)=\frac{\varphi(m)}{m},$
which yields the support claim. ∎
We shall also use the following standard Siegel–Walfisz (see [12, 13]) estimate in a range of polylogarithmic moduli.
Theorem 5 (Siegel–Walfisz).
For every fixed $A>0$ there exists $c_{A}>0$ such that uniformly for integers $m\leq(\log x)^{A}$ and reduced residue classes $a\pmod{m}$,
$\pi(x;m,a)=\frac{\operatorname{Li}(x)}{\varphi(m)}+O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr),$
where
$\pi(x;m,a)=\#\{p\leq x:p\text{ prime},\ p\equiv a\pmod{m}\}$.
For an independent, self-contained proof (based on Siegel–Walfisz) that intervals near the endpoint $1/2$ carry positive limiting measure, see Lemma 10 in Appendix A. That result is not used in what follows.
Lemma 4 (Density of even totient ratios).
The set
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}$
is dense in $\left[0,\tfrac{1}{2}\right]$.
Proof.
Since $\varphi(m)/m$ depends only on the distinct prime divisors of $m$, we have
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{p\in S}\left(1-\frac{1}{p}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
Enumerate the odd primes as $r_{1}<r_{2}<\cdots$, and set
$a_{j}:=\log\frac{r_{j}}{r_{j}-1}.$
Then $a_{j}>0$, $a_{j}\downarrow 0$, and
$a_{j}=\frac{1}{r_{j}}+O\!\left(\frac{1}{r_{j}^{2}}\right).$
Since $\sum_{j}1/r_{j}$ over the odd primes diverges, it follows that
$\sum_{j=1}^{\infty}a_{j}=+\infty.$
We claim that the set of finite subsums
$\left\{\sum_{j\in J}a_{j}:J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $0,\infty)$. Indeed, fix $y\geq 0$ and $\varepsilon>0$. Choose $J_{0}$ so large that $a_{j}<\varepsilon$ for all $j\geq J_{0}$. Since $\sum_{j\geq J_{0}}a_{j}=\infty$, there exists $N\geq J_{0}$ minimal such that
$\sum_{j=J_{0}}^{N}a_{j}>y.$
By minimality,
$\sum_{j=J_{0}}^{N-1}a_{j}\leq y<\sum_{j=J_{0}}^{N}a_{j}=\sum_{j=J_{0}}^{N-1}a_{j}+a_{N}.$
Because $a_{N}<\varepsilon$, it follows that
$y-\varepsilon<\sum_{j=J_{0}}^{N-1}a_{j}\leq y.$
Thus a finite subsum lies within $\varepsilon$ of $y$, proving the claim.
Exponentiating the negative of these finite subsums therefore shows that
$\left\{\frac{1}{2}\exp\!\left(-\sum_{j\in J}a_{j}\right):J\subset\mathbb{N}\ \text{finite}\right\}=\left\{\frac{1}{2}\prod_{j\in J}\left(1-\frac{1}{r_{j}}\right):J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
Theorem 6 (Continuous limiting distribution and full support of the totient ratio over primes).
There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
The limiting law has support exactly $0,\tfrac{1}{2}]$, with $G(\alpha)=0$ for $\alpha\leq 0$ and $G(\alpha)=1$ for $\alpha\geq 1/2$. In particular, $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
$\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
so the distribution function $G_{2}$ furnished by that proposition is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. In particular, $G_{2}$ coincides with the distribution function denoted $F$ in [4, Theorem 1.1]. Writing $G:=G_{2}$, there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
For every odd prime $p$, the integer p1p-1 is even, so
$0<\frac{\varphi(p-1)}{p-1}\leq\frac{1}{2}.$
The exceptional prime $p=2$ is negligible in the limit. Hence the limiting law is carried by $0,\tfrac{1}{2}]$, and therefore
$G(\alpha)=0\qquad(\alpha\leq 0),\qquad G(\alpha)=1\qquad(\alpha\geq 1/2).$
Let $\mu_{\varphi}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
$x_{m}:=\frac{\varphi(m)}{m}.$
By Deshouillers and Hassani [4, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
By Lemma 4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{\varphi})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{\varphi})$, every neighbourhood of $x$ has positive $\mu_{\varphi}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{\varphi}((a,b])\geq\mu_{\varphi}((x-r,x+r))>0.$
Hence
$G(b)-G(a)=\mu_{\varphi}((a,b])>0,$
so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
Theorem 6 completely determines the shifted-prime support: it is exactly $0,\tfrac{1}{2}]$. This does not settle the classical support question for the distribution of $\varphi(n)/n$ over the integers, which remains open; see [4]. The residue-class laws in [4, Proposition 2.1 and equation (2.1)] are consistent with this picture, since for primes $p\equiv 1\pmod{m}$ the corresponding limiting measure is supported in $0,\varphi(m)/m]$.
5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
The following transport lemma is the mechanism for contribution (ii) listed at the beginning of this section.
Lemma 5 (Transport lemma).
Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
If $a(p)\to 1$ as $p\to\infty$ through the primes, then $a(p)X(p)$ has the same limiting distribution $G$.
Proof.
For $\alpha<0$ there is nothing to prove, since both distribution functions vanish. Fix $\alpha\geq 0$ and $0<\varepsilon<1$. Since $a(p)\to 1$, there exists $P_{0}=P_{0}(\varepsilon)$ such that
$1-\varepsilon<a(p)<1+\varepsilon\qquad(p>P_{0},\ p\text{ prime}).$
For such primes $p$,
$X(p)\leq\frac{\alpha}{1+\varepsilon}\;\Longrightarrow\;a(p)X(p)\leq\alpha\;\Longrightarrow\;X(p)\leq\frac{\alpha}{1-\varepsilon}.$
Hence
$\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1+\varepsilon}\right\}\leq\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}+\frac{\pi(P_{0})}{\pi(x)},$
and
$\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\ a(p)X(p)\leq\alpha\bigr\}\leq\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime},\ X(p)\leq\frac{\alpha}{1-\varepsilon}\right\}+\frac{\pi(P_{0})}{\pi(x)}.$
Taking $\liminf$ and $\limsup$ as $x\to\infty$, and then letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 7 (Transfer of the limiting distribution to the PRIM-LWE density).
The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\text{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
where $G$ is the distribution function from Theorem 6.
Proof.
Write
$X(p):=\frac{\varphi(p-1)}{p-1},\qquad P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
so that $c(p)=X(p)P(p)$.
For every prime $p\geq 2$,
[…]
hence
$P(p)=1+O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 6, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
Lemma 6 (Shifted-prime pure-type and continuity criterion).
Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes. Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure. Moreover, $\mu_{f}$ is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
Proof.
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p. 258]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [3, formula (3), p. 258], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [3, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. It therefore suffices to apply the Jessen–Wintner theorem [16] (see also Tenenbaum [13, Theorem 2.7(b)]) to the infinite convolution over the odd primes. Each odd-prime factor is the characteristic function of a probability measure supported on a countable set, so the Jessen–Wintner theorem gives that the resulting infinite convolution is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of the factor at $\ell$. Since $f$ is strongly additive, the local factor at an odd prime $\ell$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. When $f(\ell)\neq 0$, these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
Lemma 7 (Prime density for a prescribed smooth kernel).
Let $y\geq 3$, and for each prime $p$ define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell$. Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. Then the set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
$\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
among the primes, in the sense that
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
Moreover, writing $C_{y}:=\prod_{3\leq\ell\leq y}(\ell-2)/(\ell-1)\asymp 1/\!\log y$, we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
Proof.
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
Let $f(n):=\log(n/\varphi(n))$, and for each prime $p$ and parameter $y\geq 3$ define
$T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then, for each fixed $y\geq 3$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$. In particular, $\eta(y)\to 0$ as $y\to\infty$.
Proof.
Since $f(\ell)=1/\ell+O(1/\ell^{2})$, we have $\eta(y)=\sum_{\ell>y}1/(\ell(\ell-1))+O(\sum_{\ell>y}1/\ell^{3})\ll 1/y$. Interchanging sums gives
$\sum_{p\leq x}T_{y}(p)\;=\;\sum_{\begin{subarray}{c}y<\ell<x\\ \ell\text{ prime}\end{subarray}}f(\ell)\,\pi(x;\ell,1).$
Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5),
$\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
Since $f(\ell)\leq 1$ for every prime $\ell$ and $\{\text{primes }\ell\in(y,L]\}\leq L$, the total error is
$\sum_{y<\ell\leq L}f(\ell)\,|E(x;\ell,1)|\;\leq\;L\cdot O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr)\;=\;o\bigl(\pi(x)\bigr).$
The main term from this range is $\operatorname{Li}(x)\sum_{y<\ell\leq L}f(\ell)/(\ell-1)=\operatorname{Li}(x)\bigl(\eta(y)+O(1/L)\bigr)$, since the tail $\sum_{\ell>L}f(\ell)/(\ell-1)\ll 1/L$.
Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let
$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
Proof.
We adapt the integer-case argument of Erdős [5] (see also [13, Exercise 256] with solution in [17, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 5) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp. 257–258] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
Step 1: Tail estimate. Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement. Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$. Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8. By that lemma and Markov’s inequality,
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (1)
which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$). Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$, for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (2)
By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (2),
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$ (3)
Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$. In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
Define the closed set
$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$. For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (1),
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
Hence, for all sufficiently large $y$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
[…]
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$. Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular. ∎
Proposition 3 (Shifted-prime limiting law for $f$).
Let
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then the empirical measures
$\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
converge weakly, as $x\to\infty$, to a Borel probability measure $\mu_{f}$ on $0,\infty)$ that is continuous and purely singular with respect to Lebesgue measure.
Proof.
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp. 257–258] for the general shift including p1p-1) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
Corollary 2 (Structure of the limiting PRIM-LWE law).
The density $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Write
$Y(p):=\frac{\varphi(p-1)}{p-1}.$
By Theorem 7, $c(p)$ has the same limiting law as $Y(p)$. By Theorem 6, this common law is continuous, has support exactly $0,\tfrac{1}{2}]$, and its distribution function is strictly increasing on $0,\tfrac{1}{2}]$. It remains only to justify pure singularity.
Set
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1}.$
Then $f$ is strongly additive and
$f(\ell)=\log\frac{\ell}{\ell-1}\sim\frac{1}{\ell}\qquad(\ell\to\infty,\ \ell\ \text{prime}),$
while
$Y(p)=e^{-f(p-1)}.$
By Proposition 3, the empirical measures of f(p1)f(p-1) converge weakly to a continuous purely singular law $\mu_{f}$. Since $T(u):=e^{-u}$ is continuous on $0,\infty)$, the continuous mapping theorem implies that
Y(p)=T(f(p1))Y(p)=T(f(p-1))
has limiting law $T_{\#}\mu_{f}$.
Choose a Borel set $E\subset[0,\infty)$ with Lebesgue measure zero and $\mu_{f}(E)=1$. The map $T$ is strictly monotone and Lipschitz on each compact interval $[m,m+1]$, $m\geq 0$. Writing $E=\bigcup_{m\geq 0}(E\cap[m,m+1])$, each image $T(E\cap[m,m+1])$ has Lebesgue measure zero (since a Lipschitz map on a compact interval preserves null sets), and therefore $T(E)$ has Lebesgue measure zero by countable subadditivity. Therefore $T_{\#}\mu_{f}$ is supported on the null set $T(E)$ and is therefore purely singular.
By Theorem 7, the same is true for $c(p)$. ∎
Note that $c(p)<1/2$ for every fixed prime $p$. Indeed, for $p=2$ one has
$\frac{\varphi(p-1)}{p-1}=\frac{\varphi(1)}{1}=1,$
so
$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)=\left(\frac{1}{2};\frac{1}{2}\right)_{\infty}\approx 0.288788<\frac{1}{2}.$
For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
Corollary 3 (Unconditional limsup).
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
Moreover, for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)$ among the primes.
Proof.
Since $c(p)<1/2$ for every prime, $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:c(p)>1/2-\varepsilon\bigr\}=1-G(1/2-\varepsilon)>0.$
In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
Corollary 4 (Small-density moduli).
For every fixed $\tau\in(0,1/2)$, both sets $\{p:c(p)\leq\tau\}$ and $\{p:c(p)>\tau\}$ have positive relative density among the primes, equal to $G(\tau)$ and $1-G(\tau)$, respectively.
Proof.
By Theorem 7 with $\alpha=\tau$, the limiting density of $\{p:c(p)\leq\tau\}$ is $G(\tau)$. Since $G$ is strictly increasing on $0,\tfrac{1}{2}]$ by Theorem 6, we have $0<G(\tau)<1$, giving positive density for both sets. ∎
Corollary 5 (Fixed-dimension analogue).
For each fixed $n\geq 1$, the exact density $c_{n}(p)$ has the same continuous limiting distribution as $\varphi(p-1)/(p-1)$; in particular, its limiting law has support exactly $0,\tfrac{1}{2}]$. Moreover,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
Write
$c_{n}(p)=\frac{\varphi(p-1)}{p-1}\,P_{n}(p),\qquad P_{n}(p):=\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Since $P_{n}(p)\to 1$ as $p\to\infty$ through the primes, Lemma 5 shows that $c_{n}(p)$ has the same limiting law as $\varphi(p-1)/(p-1)$.
Since $c_{n}(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. Let
$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
For each odd prime $p\leq x$, we have
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
so $\omega(p-1)\leq m(x)$. Let
$r_{1}<r_{2}<\cdots<r_{d}$
be the distinct prime divisors of p1p-1, where $d=\omega(p-1)$. Then $d\leq m(x)$, and $r_{i}\geq p_{i}$ for $1\leq i\leq d$. Hence
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{d}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{d}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
because each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$. Also, by the same elementary estimate used in Lemma 1,
$P_{n}(p)\geq 1-\sum_{j=1}^{n}\frac{1}{p^{j}}\geq 1-\frac{1}{p-1}\geq\frac{1}{2}.$
Therefore
$c_{n}(p)\geq\frac{1}{2}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Now $N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}p_{m(x)+1}$, so $\log x\sim\log N_{m(x)}$. Since
$\log N_{m(x)}=\sum_{i=1}^{m(x)}\log p_{i}\sim p_{m(x)},$
we get $p_{m(x)}\sim\log x$. By Mertens’ theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x},$
and therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\gg\frac{1}{\log\log x}.$
For the upper bound, the argument of Proposition 1 applies verbatim, with the additional observation that $P_{n}(p)\leq 1$ for every prime $p$. If $p^{*}(k)\leq x$ is the least prime congruent to $1\pmod{N_{k}}$ arising there, then
$c_{n}\bigl(p^{*}(k)\bigr)=\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\,P_{n}\bigl(p^{*}(k)\bigr)\leq\frac{\varphi(p^{*}(k)-1)}{p^{*}(k)-1}\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\ll\frac{1}{\log\log x}.$
Hence
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\ll\frac{1}{\log\log x},$
which proves the claim. ∎
6 Explicit Lower Bounds for Primes of Cryptographic Interest
We now give explicit lower bounds on $c(q)$ for primes $q$ whose predecessor q1q-1 has few distinct prime factors—a condition satisfied by many primes commonly used in lattice-based cryptography, especially Number Theoretic Transform (NTT)-friendly ones.
Theorem 8 (Explicit lower bound on $c(q)$ via the number of distinct prime factors).
Let $q\geq 3$ be a prime, and let $\omega=\omega(q-1)$ denote the number of distinct prime factors of q1q-1. Then
$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
Proof.
Write the distinct prime factors of q1q-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(q-1)}{q-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
since the product is minimized by choosing the smallest possible distinct primes.
Now define
$P(t):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{t^{j}}\Bigr)\qquad(t>1).$
Each factor $1-t^{-j}$ is increasing in $t$, so $P(t)$ is increasing on $(1,\infty)$. Since $q\geq 3$ is prime, we have $P(q)\geq P(3)=P_{3}$. Therefore
$c(q)=\frac{\varphi(q-1)}{q-1}\,P(q)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
as claimed. ∎
Corollary 6 (Matching lower bound).
There exists an absolute constant $c>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;\frac{c}{\log\log x}.$
Consequently,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 5: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 8 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
Corollary 7 (Pointwise universal lower bound).
Let
$m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
Then for every prime $q\geq 3$,
$c(q)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $q$,
$c(q)\geq\frac{c_{0}}{\log\log q}.$
Proof.
By Theorem 8,
$c(q)\geq P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Also,
$N_{\omega(q-1)}\leq\prod_{r\mid(q-1)}r\leq q-1,$
so $\omega(q-1)\leq m(q-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
which proves the first claim.
Finally, by Mertens’ theorem and Lemma 2,
$\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(q-1)}}\asymp\frac{1}{\log\log q},$
as $q\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(q)\ll\log\log q$. ∎
Remark 1 (On effectivity).
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ should be capable of being made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [18], Dusart [19]) with an explicit Linnik-type bound (see Xylouris [20]). We do not pursue the bookkeeping here.
Remark 2.
The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large NTT primes with $q>2^{30}$, this factor is numerically indistinguishable from 11. Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [21, 22], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
We now instantiate the general $\omega(q-1)$-based bound in several parameter regimes relevant to cryptographic applications, including many NTT-friendly moduli.
Proposition 4 (Bounds for primes with bounded $\omega(q-1)$).
Let $q\geq 3$ be a prime. Define
$B(\omega)=P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Then $c(q)\geq B(\omega(q-1))$. Explicit values:
  1. (a)
    $\omega(q-1)=1$: $B(1)=P_{3}/2\approx 0.280063$.
  2. (b)
    $\omega(q-1)\leq 5$: $B(5)\approx 0.116390$.
  3. (c)
    $\omega(q-1)\leq 10$: $B(10)\approx 0.0884704$.
  4. (d)
    $\omega(q-1)\leq 15$: $B(15)\approx 0.0776918$.
  5. (e)
    $\omega(q-1)\leq 20$: $B(20)\approx 0.0715828$.
Proof.
The bound $c(q)\geq B(\omega(q-1))$ is immediate from Theorem 8. Direct evaluation gives
$B(1)=\frac{P_{3}}{2}\approx 0.280063,\qquad B(5)=\frac{16P_{3}}{77}\approx 0.116390,$
$B(10)\approx 0.0884704,\qquad B(15)\approx 0.0776918,\qquad B(20)\approx 0.0715828.$
This yields the stated bounds. ∎
Remark 3 (On the case $\omega(q-1)=1$).
If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$. Equivalently, $q$ is a Fermat prime. In this case
$\frac{\varphi(q-1)}{q-1}=\frac{1}{2},\qquad c(q)=\frac{1}{2}\,P(q),\qquad P(q):=\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr).$
Hence $c(q)$ is close to $1/2$ for large Fermat primes; so if infinitely many Fermat primes exist, then $c(q)\to 1/2$ along that subsequence. Thus the universal bound $B(1)=P_{3}/2\approx 0.280063$ is still not tight in this special case.
Corollary 8 (Practical bounds for cryptographic primes).
Let $q$ be a prime with $q>2^{30}$. Then
$c(q)>\frac{0.56}{\log q}.$
Moreover, if $\omega(q-1)\leq 10$, then
$c(q)>0.088.$
In particular, this applies to cryptographic-size primes $q\in[2^{64},2^{128}]$ satisfying $\omega(q-1)\leq 10$.
Proof.
The claim under $\omega(q-1)\leq 10$ follows from Proposition 4(c), since
$B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
For the general bound, since every prime divisor of q1q-1 is at most q1q-1,
$\frac{\varphi(q-1)}{q-1}=\prod_{r\mid(q-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq q-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr),$
where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
By Dusart’s explicit product estimate [19, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
Also,
$\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{q^{j}}=1-\frac{1}{q-1}>0.9999.$
Therefore
$c(q)=\frac{\varphi(q-1)}{q-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(q-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(q-1)}\;>\;\frac{0.56}{\log q},$
using $e^{-\gamma}>0.5614$, $\log(q-1)<\log q$, and $0.2/\log^{3}(q-1)<2.5\times 10^{-5}$ for $q>2^{30}$. ∎
The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 7 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
Remark 4 (Typical size of $\omega(q-1)$).
Let $\Phi$ denote the standard normal distribution function. The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [23]; see also Halberstam [24] and Elliott [25, Ch. 12]) asserts that for every real $z$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\frac{\omega(q-1)-\log\log x}{\sqrt{\log\log x}}\leq z\right\}=\Phi(z).$
Consequently, for every fixed $\delta>0$,
$\frac{1}{\pi(x)}\#\left\{q\leq x:q\text{ prime and }\omega(q-1)>(1+\delta)\log\log x\right\}\longrightarrow 0.$
Thus, for large $x$, almost all primes $q\leq x$ satisfy
$\omega(q-1)\leq(1+\delta)\log\log x,$
which explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges.
Scheme $q$ factorization of q1q-1 general lower bound for $c(q)$ exact-factorization lower bound
ML-KEM $3329$ $2^{8}\cdot 13$ $B(2)=P_{3}/3\approx 0.1867$ $\dfrac{6}{13}\Bigl(1-\dfrac{1}{3328}\Bigr)\approx 0.4614$
ML-DSA $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ $B(4)=8P_{3}/35\approx 0.1280$ $\dfrac{100}{341}\Bigl(1-\dfrac{1}{8380416}\Bigr)\approx 0.2933$
Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [21] and ML-DSA [22]. The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 8, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$. The exact-factorization bounds imply expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [21, 22], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds. The exact factorizations give $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$. Together with $P(q)\geq 1-\frac{1}{q-1}$, one obtains
$c(3329)\geq\frac{6}{13}\Bigl(1-\frac{1}{3328}\Bigr)\approx 0.4614,\qquad c(8380417)\geq\frac{100}{341}\Bigl(1-\frac{1}{8380416}\Bigr)\approx 0.2933.$
A short tail-bound argument (taking $J=2$ partial products of $P(q)$ and using the estimate $P_{J}(q)-P(q)\leq q^{-J}/(q-1)$) confirms that these lower bounds are tight to many decimal places:
$c(3329)=0.4613997781\ldots,\qquad c(8380417)=0.2932550969\ldots.$
The corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.17$ and $3.41$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
Context $q$ factorization of q1q-1 $\omega(q-1)$ $1/c(q)$
ML-KEM (FIPS 203) $3329$ $^{8}\cdot 13 22$ $2.17$
ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
NewHope / FALCON [^1] FALCON is being standardized by NIST as FN-DSA (FIPS 206). [26, 27] $12289$ $^{12}\cdot 3 22$ $3.00$
BFV/BGV FHE (small) [28, 29, 30] $786433$ $^{18}\cdot 3 22$ $3.00$
SEAL BFV default [31] $132120577$ $2^{21}\cdot 3^{2}\cdot 7$ 33 $3.50$
BFV/BGV FHE (medium) [28, 29, 30] $2013265921$ $2^{27}\cdot 3\cdot 5$ 33 $3.75$
NTTRU / BLISS [32, 33] $7681$ $2^{9}\cdot 3\cdot 5$ 33 $3.75$
64-bit NTT prime [34] $2^{64}-2^{32}+1$ $2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ $2^{16}$ 11 $2.00$
Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$. All overheads are at most $4.00$. The first two rows reproduce the NIST entries from Table 1. The SEAL entry is the default coefficient modulus for poly_modulus_degree = 1024 at the 128-bit security level.
Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations. Every modulus in the table satisfies $\omega(q-1)\leq 6$ and incurs a dimension-uniform rejection-sampling overhead $1/c(q)\leq 4.00$. The table substantiates the claim that the primes typically encountered in practice have controlled factorization structure, which yields a good lower bound on $c(q)$.
Remark 5 (Concrete example).
For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [35]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [34], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$. Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so roughly a quarter of random $n\times n$ matrices have a primitive-root determinant and the security reduction loses a factor of about $4$.
7 Implications for PRIM-LWE
For a fixed prime modulus $p$, [1, Theorem 10] shows that decision-prim-lwe is at least as hard as decision-LWE, up to an $O(n^{2})$ factor. Separately, [1, Proposition 5] isolates a positive lower bound, uniform in the matrix dimension, for the probability that a random secret matrix has primitive-root determinant. In our notation, that dimension-uniform sampling constant is
$c=\inf_{n\geq 1}c_{n}(p).$
As shown in Section 2, $c_{n}(p)$ is decreasing in $n$ and $\inf_{n\geq 1}c_{n}(p)=c(p)$, so $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
Theorem 3 shows that no uniform lower bound on $c(p)$ holds across all primes, meaning the reduction tightness degrades for primes $p$ whose predecessor p1p-1 is divisible by many small primes. More precisely, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, as in the proof of Proposition 1. Then, for the absolute constants $C_{0},L$ from Theorem 4,
$N_{k}<p^{*}(k)\leq C_{0}N_{k}^{\,L}.$
The left inequality holds because $p^{*}(k)=1+jN_{k}$ with $j\geq 1$, while the right inequality is exactly Linnik’s theorem. Therefore
$\log p^{*}(k)\asymp\log N_{k},\qquad\log N_{k}\sim p_{k}$
by Lemma 2, and hence
$\log\log p^{*}(k)\asymp\log p_{k}.$
On the other hand, the proof of Proposition 1 gives
$c\bigl(p^{*}(k)\bigr)\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{k}}.$
Thus
$c\bigl(p^{*}(k)\bigr)\ll\frac{1}{\log\log p^{*}(k)},$
so the construction yields an infinite sequence of primes $p=p^{*}(k)$ for which
$c(p)=O\!\left(\frac{1}{\log\log p}\right).$
The dimension-uniform upper bound on the rejection-sampling overhead, namely $1/c(p)$, can therefore grow like
$\Omega(\log\log p).$
In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$. By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{n\geq 1}c_{n}(p)=c(p)$ is the dimension-uniform sampling constant. Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
However, this is a modulus-dependent nuance rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
Corollary 3 gives $\limsup c(p)=1/2$ unconditionally, so there is no shortage of primes for which $c(p)$ is close to its maximal value. As a concrete illustration, for safe primes $q=2r+1$ with $r$ an odd prime and $q\geq 7$, one has
$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right)\geq P_{3}\,\frac{r-1}{2r}.$
The prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$, so $c(q)$ approaches $1/2$ for large safe primes.
More generally, Corollary 4 shows that for every fixed $\tau\in(0,1/2)$, both sets of primes
$\{p:c(p)\leq\tau\}\qquad\text{and}\qquad\{p:c(p)>\tau\}$
have positive relative density, namely $G(\tau)$ and $1-G(\tau)$. Moreover, $G(\tau)\to 0$ as $\tau\downarrow 0$. Thus moduli yielding a very loose constant-factor sampling term are rare only in the threshold-asymptotic sense $\tau\downarrow 0$, not absent for any fixed threshold $\tau<1/2$.
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility. Concretely, this typically means $q\equiv 1\pmod{2^{t}}$ for moderate $t$. Mathematically, this condition controls only the 2-adic part of q1q-1. It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$. It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
Affine determinant programs (ADPs) [36] provide a natural external determinant-based connection to prim-lwe. In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial. Because the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$ rather than all of $M_{n}(\mathbb{F}_{q})$, the relevant primitive-root fraction reduces to $\varphi(q-1)/(q-1)$: the matrix-count Euler product $\prod_{j=1}^{n}(1-q^{-j})$ is already absorbed into the $\mathrm{GL}_{n}$ normalization. Since the determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{q})\longrightarrow\mathbb{F}_{q}^{\times}$
is a surjective homomorphism, all fibers have the same cardinality, and therefore the exact fraction of invertible matrices with primitive-root determinant is
$\frac{\varphi(q-1)}{q-1}.$
Hence, if $\omega(q-1)\leq 10$, then
$\frac{\varphi(q-1)}{q-1}\;\geq\;\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.15795,$
so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average. If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$. While our results do not establish a security advantage for imposing primitive-root determinant constraints, they confirm that the associated sampling overhead is mild whenever q1q-1 has controlled factorization.
8 Concluding Remarks
We have resolved the open question posed in [1, Remark 3(ii)] in the affirmative: $\inf_{p\text{ prime}}c(p)=0$, and in fact
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\text{ prime}\end{subarray}}c(p)=0.$
The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula. We have further established several results that illuminate the finer structure of the density $c(p)$:
  1. 1.
    The minimum of $c(p)$ up to $x$ has the sharp order
    $\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\asymp\frac{1}{\log\log x}.$
    In particular, the decay to zero is extremely slow and is completely quantified.
  2. 2.
    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$. This full-support conclusion appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, as shown by a transport lemma, and is purely singular. Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p. 263] noted that it carries over to shifted primes.
  3. 3.
    For the implementation moduli surveyed in Table 2—including the current NIST-standardized moduli and several widely used FHE primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$. Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
On the cryptographic side, the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but as discussed in Section 7, this represents a quantitative sensitivity to modulus selection rather than a structural obstruction. Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 7 gives $1/c(q)=O(\log\log q)$.
A natural remaining question is to characterize the limiting law $G$ more precisely—for instance, to understand its moments, its modulus of continuity, or its finer local regularity inside $0,\tfrac{1}{2}]$. This should not be confused with the classical support question for the distribution of $\varphi(n)/n$ over all integers, which remains open.
Appendix A Positive Density Near $1/2$
The following lemma gives an independent, Siegel–Walfisz-based proof that intervals near the endpoint $1/2$ carry positive limiting measure for $\varphi(p-1)/(p-1)$.
Lemma 10 (Positive density of primes with totient ratio near 1/2).
For every $\varepsilon>0$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}>0.$
Proof.
If $\varepsilon\geq 1$, then $\tfrac{1}{2}-\varepsilon\leq 0$, so the displayed set contains every prime and the claim is trivial. Hence assume throughout that $0<\varepsilon<1$.
Step 1: Sieving out small odd prime divisors. Let $R\geq 3$ be a fixed integer, to be chosen later in terms of $\varepsilon$, and define
$M_{R}:=2\!\!\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}r,\qquad\nu_{R}:=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}(r-2),\qquad\delta(R):=\frac{\nu_{R}}{\varphi(M_{R})}=\prod_{\begin{subarray}{c}3\leq r\leq R\\ r\text{ prime}\end{subarray}}\frac{r-2}{r-1}.$
Also set
[…]
The finitely many odd primes $p\leq R$ contribute only $O_{R}(1)$ to $|\mathcal{A}_{R}(x)|$. Hence, by Theorem 5,
$|\mathcal{A}_{R}(x)|=\frac{\nu_{R}}{\varphi(M_{R})}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\delta(R)\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr).$
Applying Theorem 5 also with modulus $1$, we have
$\pi(x)=\operatorname{Li}(x)+O\!\bigl(xe^{-c\sqrt{\log x}}\bigr),$
and therefore
$\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}\longrightarrow\delta(R)>0.$
In particular, $\mathcal{A}_{R}(x)$ has positive relative density $\delta(R)$ among the primes.
Step 2: Reduction to a mean-value estimate. For $p\in\mathcal{A}_{R}(x)$, the only prime $\leq R$ dividing p1p-1 is $2$, so
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2}\prod_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\Bigl(1-\frac{1}{q}\Bigr).$
Set
$S(p):=\sum_{\begin{subarray}{c}q\mid(p-1)\\ q>R\end{subarray}}\frac{1}{q}.$
By the elementary inequality
$\prod_{i}(1-a_{i})\geq 1-\sum_{i}a_{i}\qquad(0\leq a_{i}\leq 1),$
any $p\in\mathcal{A}_{R}(x)$ with $S(p)\leq\varepsilon$ satisfies
$\frac{\varphi(p-1)}{p-1}\geq\frac{1}{2}\bigl(1-S(p)\bigr)\geq\frac{1}{2}(1-\varepsilon)>\frac{1}{2}-\varepsilon.$
It therefore suffices to show that a positive proportion of primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$.
Step 3: Mean-value estimate for $S(p)$. Throughout Steps 3 and 4, the parameter $R$ (and hence $\varepsilon$) is fixed; all limits are as $x\to\infty$. Fix $Q=(\log x)^{2}$, and decompose $S(p)=S_{\leq Q}(p)+S_{>Q}(p)$, where
$S_{\leq Q}(p)=\sum_{\begin{subarray}{c}q\mid(p-1)\\ R<q\leq Q\end{subarray}}\frac{1}{q}.$
For the tail, each prime divisor $q>Q$ of p1p-1 contributes at most $1/Q$, and there are at most $\log p/\log Q$ such divisors. Hence, uniformly for $p\leq x$,
$S_{>Q}(p)\leq\frac{\log p}{Q\log Q}\leq\frac{\log x}{Q\log Q}=\frac{1}{2\log x\,\log\log x}=o(1).$
Now assume $x$ is large enough that $Q>R$. For each prime $q\in(R,Q]$, write
$B_{q}(x):=\#\{p\in\mathcal{A}_{R}(x):p\equiv 1\pmod{q}\}$.
Since $q>R$, we have $q\nmid M_{R}$. The conditions $p\in\mathcal{A}_{R}(x)$ and $p\equiv 1\pmod{q}$ therefore define exactly $\nu_{R}$ reduced residue classes modulo $M_{R}q$: modulo each odd prime $r\leq R$ there are r2r-2 admissible reduced classes, modulo $2$ there is exactly a single admissible class, and modulo $q$ the residue class is forced to be $1$. Because $\varphi(M_{R}q)=\varphi(M_{R})(q-1)$, we obtain from Theorem 5 that
$B_{q}(x)=\frac{\nu_{R}}{\varphi(M_{R}q)}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr)=\frac{\delta(R)}{q-1}\,\operatorname{Li}(x)+O_{R}\!\bigl(xe^{-c_{R}\sqrt{\log x}}\bigr),$
uniformly for $R<q\leq Q$. Here we use that $M_{R}$ is fixed and $q\leq(\log x)^{2}$, so $M_{R}q\leq(\log x)^{A}$ for some fixed $A$ and all sufficiently large $x$.
Dividing by the asymptotic formula for $|\mathcal{A}_{R}(x)|$ yields
$\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}=\frac{1}{q-1}+O_{R}\!\bigl((\log x)e^{-c_{R}\sqrt{\log x}}\bigr)$
uniformly for $R<q\leq Q$. Therefore
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S_{\leq Q}(p)$ $=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q}\,\frac{B_{q}(x)}{|\mathcal{A}_{R}(x)|}$
$\leq\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+O_{R}\!\Bigl((\log x)e^{-c_{R}\sqrt{\log x}}\sum_{q\leq Q}\frac{1}{q}\Bigr)$
$=\sum_{\begin{subarray}{c}R<q\leq Q\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}+o(1),$
since $\sum_{q\leq Q}1/q\ll\log\log Q\ll\log\log\log x$.
Combining this with the uniform bound on $S_{>Q}(p)$, we get
$\frac{1}{|\mathcal{A}_{R}(x)|}\sum_{p\in\mathcal{A}_{R}(x)}S(p)\leq\eta(R)+o(1),\qquad\eta(R):=\sum_{\begin{subarray}{c}q>R\\ q\text{ prime}\end{subarray}}\frac{1}{q(q-1)}.$
Since $\sum_{q}1/(q(q-1))$ converges, we have $\eta(R)\to 0$ as $R\to\infty$.
Step 4: Markov’s inequality and conclusion. Choose $R$ large enough that $\eta(R)<\varepsilon/2$. By Markov’s inequality, the fraction of primes in $\mathcal{A}_{R}(x)$ with $S(p)>\varepsilon$ is at most
$\frac{\eta(R)+o(1)}{\varepsilon}<\frac{1}{2}$
for all sufficiently large $x$. Hence at least half the primes in $\mathcal{A}_{R}(x)$ satisfy $S(p)\leq\varepsilon$, and therefore
$\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon.$
It follows that for all sufficiently large $x$,
$\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}\geq\frac{1}{2}\,|\mathcal{A}_{R}(x)|.$
Dividing by $\pi(x)$ and taking $\liminf$, we conclude that
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\left\{p\leq x:p\text{ prime and }\frac{\varphi(p-1)}{p-1}>\frac{1}{2}-\varepsilon\right\}$ $\geq\frac{1}{2}\lim_{x\to\infty}\frac{|\mathcal{A}_{R}(x)|}{\pi(x)}$
$=\frac{\delta(R)}{2}>0.$
This proves the lemma. ∎