### 21 Apr 2026 v4

# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

Abstract
For a prime $p$, let $c_{n}(p)$ denote the fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$, and let $c(p)=\lim_{n\to\infty}c_{n}(p)$ be the limiting density. This quantity governs the efficiency of PRIM-LWE, a variant of the learning with errors (LWE) problem in which the secret matrix is required to have primitive-root determinant. The value $1/c(p)$ equals the expected rejection-sampling overhead in the reduction from standard LWE to PRIM-LWE. We prove unconditionally that $\inf_{p\text{ prime}}c(p)=0$, resolving an open question from the PRIM-LWE literature, and establish the rate $\min_{p\leq x}c(p)\asymp 1/\log\log x$ as $x\to\infty$. We show that $c(p)$, viewed as a function on the primes, has a continuous, purely singular limiting distribution with support $[0,1/2]$ and Hausdorff dimension 0. The moments of this limiting distribution are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\operatorname{Re}(s)>0$. Using the number of distinct prime factors $\omega(p-1)$, we derive explicit lower bounds: for every prime $p>2^{30}$ the overhead satisfies $1/c(p)<1.79\log p$, and in general $1/c(p)=O(\log\log p)$. We also describe a certified algorithm for evaluating $c(p)$ and tabulate values for primes of cryptographic interest.
Keywords: Primitive root, determinant density, PRIM-LWE, limiting distribution, Hausdorff dimension
MSC 2020: 11A07, 11N60, 94A60, 15B33
1 Introduction
Let $p$ be a prime and $M_{n}(\mathbb{F}_{p})$ denote the ring of $n\times n$ matrices over $\mathbb{F}_{p}$. Following [36], a matrix $M\in M_{n}(\mathbb{F}_{p})$ has primitive-root determinant if $\det(M)$ generates $\mathbb{F}_{p}^{*}$. Write $c_{n}(p)$ for the fraction of such matrices, and $c(p):=\inf_{n\geq 1}c_{n}(p)=\lim_{n\to\infty}c_{n}(p)$ for the limiting density; the precise formulas are recorded in Section 2. Since $c_{n}(p)$ is decreasing in $n$, one has $c(p)\leq c_{n}(p)$ for every $n\geq 1$. Consequently, every lower bound on $c(p)$ is simultaneously a lower bound on every finite-dimensional density $c_{n}(p)$.
The quantity $c(p)$ arises as a reduction constant in the PRIM-LWE variant [36] of the popular learning with errors (LWE) problem [33], introduced to construct verifiable secret sharing schemes with “free” verification. The reduction proceeds by rejection-sampling and succeeds with probability $c_{n}(p)$ in dimension $n$, so the dimension-uniform overhead is $1/c(p)$. This overhead represents a tightness loss in the security reduction: the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates. Sehrawat, Yeo, and Desmedt [36, Remark 3(ii)] asked whether $\inf_{p}c(p)=0$ can be proved unconditionally.
Main results
1. Unconditional vanishing (Theorem 2). We prove $\inf_{p\text{ prime}}c(p)=0$ using only Dirichlet’s theorem on primes in arithmetic progressions and Mertens’ third theorem, without assuming the primorial-prime hypothesis; see Section 3.
2. Sharp order of the minimum (Proposition 1, Corollary 7). We establish
$\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
See Remark 4 for effectivity.
3. Limiting distribution (Theorems 4 and 5, Corollary 2, Proposition 4). The quantity $c(p)$ possesses a continuous, purely singular limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the limiting measure has Hausdorff dimension 0. The full-support conclusion is new; it combines the left non-differentiability result of Deshouillers–Hassani [9, Theorem 1.1] with the density of even totient ratios (Lemma 4). The moments of $G$ are given by convergent Euler products (Theorem 6), the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 8), and $G$ is not Hölder of any positive order at $1/2$ (Proposition 7).
Since the limiting law is purely singular and supported on a set of Hausdorff dimension 0, it cannot certify $c(p)$ at any individual prime; this motivates the pointwise bounds of item 4.
4. Explicit bounds via distinct prime factors (Theorem 9). We give unconditional lower bounds on $c(p)$ for primes $p$ parameterized by $\omega(p-1)$, the number of distinct prime factors of p1p-1, yielding computable bounds for any prime $p$ such that the factorization of p1p-1 is known. For instance, for the NIST-standardized primes $p=3329$ and $p=8380417$, the exact factorizations of p1p-1 give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$; see Section 7. In particular, the dimension-uniform expected rejection-sampling overhead $1/c(p)$ in the LWE-to-PRIM-LWE reduction [36, Theorem 10] admits an explicit upper bound depending only on $\omega(p-1)$.
From the computational viewpoint, our main contributions are: explicit lower bounds on $c(p)$ valid for all primes $p>2^{30}$ (Corollary 9), a certified rational-arithmetic algorithm for evaluating $c(p)$ to arbitrary precision given the factorization of p1p-1 (Proposition 8), and tables of $c(p)$ for all primes appearing in current NIST post-quantum cryptographic standards and representative lattice-based implementations (Tables 2 and 3). The explicit bound takes the following compact form:
Corollary 1 (Cryptographic consequence).
Let $P_{3}:=\prod_{j=1}^{\infty}(1-1/3^{j})\approx 0.5601260779$. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then
$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
The general $\omega(p-1)$-based bound is immediate from Theorem 9. Table 1 lists representative values of $B(K)$, and Tables 2 and 3 collect these bounds for representative cryptographic moduli.
Related work
A related distribution problem was studied by Aivazidis and Sofos [1], who considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ as $q$ varies over prime powers and proved corresponding distribution laws. Their setting is different from ours: we work over prime fields and with the primitive-root determinant density arising in PRIM-LWE. On the distributional side, Deshouillers and Hassani [9, Proposition 2.1 and equation (2.1)] proved limiting-distribution laws for $\varphi(p-1)/(p-1)$ in the residue classes $p\equiv 1\pmod{m}$, including continuity, and [9, Theorem 1.1] established left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even. The analogous support question for the classical distribution of $\varphi(n)/n$ over the integers remains open; see [9].
2 Background and the Open Question
Throughout, $\log$ denotes the natural logarithm. We write $\varphi$ for Euler’s totient function, $\omega(n)$ for the number of distinct prime factors of $n$, and $p_{k}$ for the $k$-th prime.
Following [36], define
$M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
The determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{p})\longrightarrow\mathbb{F}_{p}^{\times}$
is a surjective group homomorphism, so a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
(see, e.g., Mullen and Panario [30, Theorem 13.3.3]), it follows that
$c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
cf. [36, Corollary 4]. Consequently, the limit
$c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
exists and is strictly positive for each fixed $p$. Since $c_{n}(p)$ is decreasing in $n$, $\inf_{n\geq 1}c_{n}(p)=c(p)$.
Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorization of p1p-1, one has
$c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$ (1)
Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behavior of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
In [36, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [7] for computational results on primorial primes.
3 Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results: Dirichlet’s theorem on primes in arithmetic progressions [10] and the following.
Theorem 1 (Mertens’ Third Theorem [28]; see [29, Theorem 2.7(e)], 1874).
As $x\to\infty$,
$\prod_{\begin{subarray}{c}\ell\leq x\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{\ell}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
Lemma 1 (Lower bound for the matrix-count Euler product).
For all primes $p\geq 3$,
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;\frac{1}{2}.$
Proof.
By the finite-product inequality $\prod(1-a_{j})\geq 1-\sum a_{j}$ (immediate by induction when $0\leq a_{j}\leq 1$) and the geometric series $\sum_{j\geq 1}p^{-j}=1/(p-1)\leq 1/2$. ∎
Theorem 2 (Unconditional vanishing of $\inf c(p)$).
The following identities hold:
$\inf_{p\textup{ prime}}c(p)=0,\qquad\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0.$
Proof.
For each $k\geq 1$ define the $k$th primorial
$N_{k}\;=\;\prod_{i=1}^{k}p_{i}.$
Since $\gcd(1,N_{k})=1$, Dirichlet’s theorem guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $\ell\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
By Mertens’ third theorem (Theorem 1),
$\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since the infinite product satisfies $\prod_{j=1}^{\infty}(1-p^{-j})\leq 1$ for every prime $p$, we obtain
$c(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since for every $k$ there are infinitely many primes $p\equiv 1\pmod{N_{k}}$, the displayed upper bound holds for arbitrarily large primes $p$. Letting $k\to\infty$ gives
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0,$
and the infimum statement follows immediately. ∎
4 Quantitative Refinements
By Linnik’s theorem [26], there exist absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
Let $N_{k}$ denote the $k$-th primorial, $\vartheta$ the Chebyshev function, and $m(x)$ the primorial-counting function:
$N_{k}:=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p,\qquad m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
Lemma 2 (Prime-number-theorem consequences for primorials; cf. [29, Corollary 2.5]).
As $k\to\infty$,
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}.$
As $x\to\infty$,
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x.$
Proof.
The first two assertions follow from $\vartheta(x)\sim x$. For the last two, $N_{m(x)}\leq x<N_{m(x)}p_{m(x)+1}$ gives $\log x\sim\log N_{m(x)}\sim p_{m(x)}$, since $\log p_{m(x)+1}=o(p_{m(x)})$. ∎
Proposition 1 (Upper bound for the minimum density up to $x$).
There exists an absolute constant $C>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;\frac{C}{\log\log x}.$
Proof.
Fix absolute constants $C_{0},L>0$ from Linnik’s theorem such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that
$C_{0}N_{k}^{\,L}\leq x.$
Then
$p^{*}(k)\leq C_{0}N_{k}^{\,L}\leq x.$
Such a $k$ exists for all sufficiently large $x$, and $k\to\infty$ as $x\to\infty$.
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x,$
and hence $p_{k}\ll\log x$. Therefore $\log p_{k}=O(\log\log x)$.
Conversely, the maximality of $k$ forces $C_{0}N_{k+1}^{\,L}>x$, so
$L\log N_{k+1}>\log x-\log C_{0}.$
Since $\log N_{k+1}=\log N_{k}+\log p_{k+1}$, while
$\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}\quad\text{and}\quad\log p_{k+1}\sim\log p_{k}=o(p_{k})=o(\log N_{k}),$
we have $\log N_{k+1}\sim\log N_{k}$. Hence $\log N_{k}\gg\log x$, so $p_{k}\gg\log x$ and therefore $\log p_{k}=\Omega(\log\log x)$. Combined with the upper bound just proved, this yields
$\log p_{k}\asymp\log\log x.$
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 2 gives
$c\bigl(p^{*}(k)\bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\asymp\;\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;c\bigl(p^{*}(k)\bigr)\;\ll\;\frac{1}{\log\log x},$
as claimed. ∎
The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7 below; its proof requires only the primorial counting argument and Lemma 1.
5 The Limiting Distribution of $c(p)$
The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [16] (Tenenbaum [38, pp. 262–263], Deshouillers–Hassani [9]). Throughout this section we write
$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (2)
5.1 The distribution of $\varphi(p-1)/(p-1)$
Lemma 3 (Verification of the Erdős–Wintner hypotheses).
The function $f$ defined in (2) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
The three convergence conditions of the shifted-prime Erdős–Wintner theorem hold:
$\sum_{|f(\ell)|>1}\frac{1}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,$
as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same conclusions hold for $g(n)=\varphi(n)/n$.
Proof.
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied. The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$. ∎
Proposition 2 (Deshouillers–Hassani; special case of [9, Proposition 2.1 and equation (2.1)]).
For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\left\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\right\}=G_{m}(\alpha).$
Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
Proof.
Apply Deshouillers and Hassani [9, Proposition 2.1] to
$g(n):=\frac{\varphi(n)}{n}.$
By Lemma 3 applied to $\log g(\ell)=-\log(\ell/(\ell-1))$, the three convergence hypotheses of [9, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the continuity criterion in the same proposition, $G_{m}$ is continuous if and only if
$\sum_{\begin{subarray}{c}\ell\equiv 1\pmod{m}\\ g(\ell)\neq 1\end{subarray}}\frac{1}{\ell}$
diverges. Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
$\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [29, Corollary 4.10]). Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
If $p\equiv 1\pmod{m}$, then every prime divisor of $m$ also divides p1p-1, so
$\frac{\varphi(p-1)}{p-1}=\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\leq\prod_{\ell\mid m}\Bigl(1-\frac{1}{\ell}\Bigr)=\frac{\varphi(m)}{m},$
which yields the support claim. ∎
We shall also use the following standard Siegel–Walfisz (see [29, Corollary 11.21] or [39, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
Theorem 3 (Siegel–Walfisz).
For every fixed $A>0$ there exists $c_{A}>0$ such that uniformly for integers $m\leq(\log x)^{A}$ and reduced residue classes $a\pmod{m}$,
$\pi(x;m,a)=\frac{\operatorname{Li}(x)}{\varphi(m)}+O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr),$
where
$\pi(x;m,a)=\#\{p\leq x:p\textup{ prime},\ p\equiv a\pmod{m}\}$.
Lemma 4 (Density of even totient ratios).
The set
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}$
is dense in $\left[0,\tfrac{1}{2}\right]$.
Proof.
Since $\varphi(m)/m$ depends only on the distinct prime divisors of $m$, we have
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{r\in S}\left(1-\frac{1}{r}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
Enumerate the odd primes as $r_{1}<r_{2}<\cdots$, and set
$a_{j}:=\log\frac{r_{j}}{r_{j}-1}.$
Then $a_{j}>0$, $a_{j}\downarrow 0$, and
$a_{j}=\frac{1}{r_{j}}+O\!\left(\frac{1}{r_{j}^{2}}\right).$
Since $\sum_{j}1/r_{j}$ over the odd primes diverges, it follows that
$\sum_{j=1}^{\infty}a_{j}=+\infty.$
We claim that the set of finite subsums
$\left\{\sum_{j\in J}a_{j}:J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $0,\infty)$. Indeed, fix $y\geq 0$ and $\varepsilon>0$. Choose $J_{0}$ so large that $a_{j}<\varepsilon$ for all $j\geq J_{0}$. Since $\sum_{j\geq J_{0}}a_{j}=\infty$, there exists $N\geq J_{0}$ minimal such that
$\sum_{j=J_{0}}^{N}a_{j}>y.$
By minimality,
$\sum_{j=J_{0}}^{N-1}a_{j}\leq y<\sum_{j=J_{0}}^{N}a_{j}=\sum_{j=J_{0}}^{N-1}a_{j}+a_{N}.$
Because $a_{N}<\varepsilon$, it follows that
$y-\varepsilon<\sum_{j=J_{0}}^{N-1}a_{j}\leq y.$
Thus a finite subsum lies within $\varepsilon$ of $y$, proving the claim.
Exponentiating the negative of these finite subsums therefore shows that
$\left\{\frac{1}{2}\exp\!\left(-\sum_{j\in J}a_{j}\right):J\subset\mathbb{N}\ \text{finite}\right\}=\left\{\frac{1}{2}\prod_{j\in J}\left(1-\frac{1}{r_{j}}\right):J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
Theorem 4 (Limiting distribution and full support over primes).
There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
The limiting law has support exactly $0,\tfrac{1}{2}]$, with $G(\alpha)=0$ for $\alpha\leq 0$ and $G(\alpha)=1$ for $\alpha\geq 1/2$. In particular, $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Apply Proposition 2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
$\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. Writing $G:=G_{2}$ (denoted $F$ in [9, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
For every odd prime $p$, the integer p1p-1 is even, so
$0<\frac{\varphi(p-1)}{p-1}\leq\frac{1}{2}.$
The exceptional prime $p=2$ is negligible in the limit. Hence the limiting law is carried by $0,\tfrac{1}{2}]$, and therefore
$G(\alpha)=0\qquad(\alpha\leq 0),\qquad G(\alpha)=1\qquad(\alpha\geq 1/2).$
Let $\mu_{\varphi}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
$x_{m}:=\frac{\varphi(m)}{m}.$
By Deshouillers and Hassani [9, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{\varphi})$, then some open interval $I$ containing $x_{m}$ has $\mu_{\varphi}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{\varphi})$ for every even $m$.
By Lemma 4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{\varphi})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{\varphi})$, every neighbourhood of $x$ has positive $\mu_{\varphi}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{\varphi}((a,b])\geq\mu_{\varphi}((x-r,x+r))>0.$
Hence
$G(b)-G(a)=\mu_{\varphi}((a,b])>0,$
so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
The support question for $\varphi(n)/n$ over the integers remains open; see Schoenberg [35] and [9].
5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 5 (Transport lemma).
Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
If $a(p)\to 1$ as $p\to\infty$ through the primes, then $a(p)X(p)$ has the same limiting distribution $G$.
Proof.
Fix $\alpha\geq 0$ and $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
$\{X(p)\leq\alpha/(1+\varepsilon)\}\;\subseteq\;\{a(p)X(p)\leq\alpha\}\;\subseteq\;\{X(p)\leq\alpha/(1-\varepsilon)\}$
up to finitely many exceptions. Taking densities as $x\to\infty$ and letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 5 (Transfer to the PRIM-LWE density).
The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
where $G$ is the distribution function from Theorem 4.
Proof.
Write
$X(p):=\frac{\varphi(p-1)}{p-1},\qquad P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
so that $c(p)=X(p)P(p)$.
For every prime $p\geq 2$,
[…]
hence
$P(p)=1+O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 4, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
5.3 Pure singularity
Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law. By the Jessen–Wintner theorem [24], any limiting distribution arising from an infinite convolution of discrete measures must be of pure type—purely discrete, purely singular, or absolutely continuous. Continuity was already verified in Lemma 3, so the remaining question is whether the law is absolutely continuous or purely singular. The latter turns out to hold: the values of $c(p)$ cluster on a fractal-like set of Lebesgue measure zero, as we prove by adapting a classical argument of Erdős to the shifted-prime setting.
Lemma 6 (Shifted-prime pure-type and continuity criterion).
Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes. Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure. Moreover, $\mu_{f}$ is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
Proof.
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [25, Theorem 2] and Hildebrand [23, Theorem 5] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem [4] (see also [39, Theorem II.8.34]): the only property of the shift used in these proofs is the Bombieri–Vinogradov estimate for primes in arithmetic progressions, which holds uniformly for any fixed nonzero $a$. Hildebrand [23, p. 211] notes explicitly that the same methods apply to $\{p+a\}$ for arbitrary nonzero $a$. For the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [38, formula (3), p. 262]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [38, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{\ell\textup{ prime}}\!\left(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\right).$
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [38, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. For each odd prime $\ell$, the remaining local factor is the characteristic function of an atomic probability measure $\nu_{\ell}$ placing mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. The product formula therefore represents the characteristic function of the infinite convolution $\delta_{f(2)}\ast\mathop{{\ast}}_{\ell\geq 3}\nu_{\ell}$. By Lévy’s uniqueness theorem (see, e.g., [39, Section III.2]), $\mu_{f}$ equals this infinite convolution. It therefore suffices to apply the Jessen–Wintner theorem (see also Tenenbaum [39, Theorem III.2.7(b)]) to the convolution over the odd primes. Each factor $\nu_{\ell}$ is supported on a countable (indeed two-element) set, so the Jessen–Wintner theorem gives that the resulting infinite convolution—and hence $\mu_{f}$ itself—is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of $\nu_{\ell}$. Since $\nu_{\ell}$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$, when $f(\ell)\neq 0$ these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
For $y\geq 3$ and each prime $p$, define
$a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y\\ \ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell,$
and write $P^{+}(n)$ for the largest prime factor of $n$.
Lemma 7 (Prime density for a prescribed smooth kernel).
Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. The set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
$\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
among the primes, in the sense that
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
Moreover, writing
$C_{y}:=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
Proof.
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 3) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem [29, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
Let $f$ be as in (2), and for each prime $p$ and parameter $y\geq 3$ define
$T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then, for each fixed $y\geq 3$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$.
Proof.
Since $f(\ell)=1/\ell+O(1/\ell^{2})$, we have $\eta(y)=\sum_{\ell>y}1/(\ell(\ell-1))+O(\sum_{\ell>y}1/\ell^{3})\ll 1/y$. Interchanging sums gives
$\sum_{p\leq x}T_{y}(p)\;=\;\sum_{\begin{subarray}{c}y<\ell<x\\ \ell\text{ prime}\end{subarray}}f(\ell)\,\pi(x;\ell,1).$
Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 3),
$\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
Since $f(\ell)\leq 1$ for every prime $\ell$ and $\{\text{primes }\ell\in(y,L]\}\leq L$, the total error is
$\sum_{y<\ell\leq L}f(\ell)\,|E(x;\ell,1)|\;\leq\;L\cdot O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr)\;=\;o\bigl(\pi(x)\bigr).$
The main term from this range is $\operatorname{Li}(x)\sum_{y<\ell\leq L}f(\ell)/(\ell-1)=\operatorname{Li}(x)\bigl(\eta(y)+O(1/L)\bigr)$, since the tail $\sum_{\ell>L}f(\ell)/(\ell-1)\ll 1/L$.
Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let $f$ be the strongly additive function defined in (2). The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
Proof.
We adapt the integer-case argument of Erdős [14, pp. 722–725] (see also [39, Exercise 256] with solution in [40, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 3) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Lemma 6), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
Step 1: Tail estimate. Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement. Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$. Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8. By that lemma and Markov’s inequality,
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (3)
which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$). Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$ (see, e.g., [39, Theorem I.3.10]), for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (4)
By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (4),
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$
Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$. In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
Define the closed set
$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$. For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (3),
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
Hence, for all sufficiently large $y$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
[…]
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$. Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular. ∎
Proposition 3 (Shifted-prime limiting law for $f$).
Let $f$ be as in (2). Then the empirical measures
$\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
converge weakly, as $x\to\infty$, to a Borel probability measure $\mu_{f}$ on $0,\infty)$ that is continuous and purely singular with respect to Lebesgue measure.
Proof.
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Lemma 6) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
Corollary 2 (Structure of the limiting PRIM-LWE law).
The density $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Combine Theorems 4 and 5: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
By Proposition 3, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (2)) is continuous and purely singular. Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is locally Lipschitz on $0,\infty)$, the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets and is therefore purely singular. By Theorem 5, the same holds for $c(p)$. ∎
Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
Figure 1: Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical limiting distribution function $G$ of Theorem 5, estimated by $10^{7}$ Monte Carlo samples of the infinite convolution of Lemma 10 truncated at the 10,00010{,}000th odd prime. The empirical values of $c(p)$ were computed via the truncated Euler product of Proposition 8 with $J=20$ terms, contributing error below $10^{-9}$ for every prime $p\geq 3$.
Remark 1.
The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$. Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.288788<1/2$. For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support of the limiting law for $c(p)$, not an attained value.
Corollary 3 (Unconditional limsup).
We have
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=\frac{1}{2}.$
Moreover, for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)$ among the primes.
Proof.
Since $c(p)<1/2$ for every prime (Remark 1), $\limsup c(p)\leq 1/2$. For the reverse inequality, fix $\varepsilon>0$. By Corollary 2, the limiting distribution of $c(p)$ has support exactly $0,\tfrac{1}{2}]$, so the limiting measure assigns positive mass to the interval $(1/2-\varepsilon,\,1/2]$. Hence
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:c(p)>1/2-\varepsilon\bigr\}=1-G(1/2-\varepsilon)>0.$
In particular, infinitely many primes satisfy $c(p)>1/2-\varepsilon$, so $\limsup c(p)\geq 1/2-\varepsilon$. Letting $\varepsilon\downarrow 0$ yields the claim. ∎
Corollary 4 (Fixed-dimension analogue).
For each fixed $n\geq 1$, the exact density $c_{n}(p)$ has the same continuous limiting distribution as $\varphi(p-1)/(p-1)$; in particular, its limiting law has support exactly $0,\tfrac{1}{2}]$. Moreover,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c_{n}(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
Write $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})$. Since $P_{n}(p)\to 1$, Lemma 5 gives the distributional claim. For the lower bound on $\min_{p\leq x}c_{n}(p)$: every odd prime $p\leq x$ satisfies $\omega(p-1)\leq m(x)$, so
$\frac{\varphi(p-1)}{p-1}\;\geq\;\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
and $P_{n}(p)\geq 1/2$ by Lemma 1. By Lemma 2 and Mertens’ third theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
The upper bound is identical to Proposition 1, since $P_{n}(p)\leq 1$. ∎
6 Hausdorff Dimension, Moments, and Fourier Structure of $G$
The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure? Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension. We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support is the full interval $0,\tfrac{1}{2}]$; in particular, the fractal-like clustering visible in Figure 2 is not an artifact of finite sampling. We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform. The section concludes with a precise asymptotic for the behavior of $G$ near the endpoint $1/2$.
Throughout, $f$ is the strongly additive function defined in (2), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$.
Lemma 10 (Factor structure of $\mu_{f}$).
The measure $\mu_{f}$ is the infinite convolution
$\mu_{f}\;=\;\delta_{\log 2}\;*\;\mathop{{\ast}}_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\nu_{\ell},$
where $\delta_{\log 2}$ is the point mass at $\log 2$ and, for each odd prime $\ell$,
$\nu_{\ell}\;:=\;\frac{\ell-2}{\ell-1}\,\delta_{0}\;+\;\frac{1}{\ell-1}\,\delta_{f(\ell)},\qquad f(\ell)=\log\frac{\ell}{\ell-1}.$
Equivalently, $Y\overset{d}{=}\log 2+\sum_{\ell\geq 3,\;\ell\textup{ prime}}Z_{\ell}$ with the $Z_{\ell}$ independent, $\mathbb{P}(Z_{\ell}=0)=(\ell-2)/(\ell-1)$, $\mathbb{P}(Z_{\ell}=f(\ell))=1/(\ell-1)$.
Proof.
By Tenenbaum [38, formula (3), p. 262] for the shift a=1a=-1, the characteristic function of $\mu_{f}$ is
$\widehat{\mu_{f}}(\tau)=\prod_{\ell\text{ prime}}\Bigl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\Bigr).$
Since $f$ is strongly additive, $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, whence $\sum_{\nu\geq 1}e^{i\tau f(\ell^{\nu})}\ell^{-\nu}=e^{i\tau f(\ell)}/(\ell-1)$. At each odd prime $\ell$ the local factor equals $\frac{\ell-2}{\ell-1}+\frac{e^{i\tau f(\ell)}}{\ell-1}=\widehat{\nu_{\ell}}(\tau)$; at $\ell=2$ it equals $e^{i\tau\log 2}=\widehat{\delta_{\log 2}}(\tau)$. Since the characteristic function determines a probability measure uniquely, the convolution decomposition follows. ∎
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [17, Chapter 2]):
$\dim_{H}(\mu):=\inf\{\dim_{H}(E):E\subset\mathbb{R}\text{ is Borel and }\mu(E)=1\}.$
Proposition 4 (Hausdorff dimension zero of the limiting law).
$\dim_{H}(\mu_{f})=0$. Consequently, if $\mu_{G}$ denotes the limiting measure on $0,\tfrac{1}{2}]$ with distribution function $G$, then $\dim_{H}(\mu_{G})=0$.
Proof.
By Lemma 10,
$Y\overset{d}{=}\log 2+\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad w_{\ell}:=\log\frac{\ell}{\ell-1},$
where the $B_{\ell}$ are independent Bernoulli random variables with
$\mathbb{P}(B_{\ell}=1)=\frac{1}{\ell-1},\qquad\mathbb{P}(B_{\ell}=0)=\frac{\ell-2}{\ell-1}.$
Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure.
For $y\geq 3$, let
$I(y):=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right).$
If $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\boldsymbol{b}\bigr)=e^{-I_{y}(\boldsymbol{b})},$
where $I_{y}(\boldsymbol{b})$ denotes the value of $I(y)$ at that configuration. Moreover,
$\mathbb{E}I(y)\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{\log\ell}{\ell}\ll\log y$
and
$\operatorname{Var}(I(y))\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{2}\right]\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\ll(\log y)^{2}.$
Hence there exists an absolute constant $C_{0}>0$ such that
$\mathbb{E}I(y)\leq C_{0}\log y,\qquad\operatorname{Var}(I(y))\leq C_{0}(\log y)^{2}\qquad(y\geq 3).$
Choose $C:=4C_{0}$ (so that $Cn-a_{n}s/2\leq-Cn$ in the covering estimate below).
$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2}.$
We construct the sequence $(y_{n})$ recursively. Set $y_{1}:=3$. Given $y_{n}$, define $a_{n}$, $z_{n}$, $r_{n}$ as above and put $y_{n+1}:=z_{n}+1$. Then the intervals $(y_{n},z_{n}]$ are pairwise disjoint by construction, $z_{n}=y_{n}^{a_{n}}$ grows super-exponentially, and the summability conditions $\sum r_{n}<\infty$ and $\sum y_{n}^{-Cn}<\infty$ are immediate since $r_{n}=z_{n}^{-1/2}$ and $y_{n}^{-Cn}\leq y_{n}^{-C}$ with $y_{n}\to\infty$ super-exponentially.
For each $n$, set $I_{n}:=I(y_{n})$ and
$A_{n}:=\{I_{n}\leq Cn\log y_{n}\}.$
By Chebyshev’s inequality,
$\mathbb{P}(A_{n}^{c})\leq\frac{\operatorname{Var}(I_{n})}{(Cn\log y_{n}-\mathbb{E}I_{n})^{2}}\ll\frac{1}{n^{2}}.$
Let $T_{n}$ be the set of prefix configurations $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\boldsymbol{b})\leq Cn\log y_{n}$. Then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}\bigr)=1-O(n^{-2}),$
and every atom in $T_{n}$ has probability at least $y_{n}^{-Cn}$. Since the total mass is at most $1$, it follows that
$|T_{n}|\leq y_{n}^{Cn}.$
Now define the empty-block event
$E_{n}:=\{B_{\ell}=0\text{ for every prime }\ell\in(y_{n},z_{n}]\}.$
Because the blocks $(y_{n},z_{n}]$ are disjoint, the events $E_{n}$ are independent. Moreover,
$\mathbb{P}(E_{n})=\prod_{\begin{subarray}{c}y_{n}<\ell\leq z_{n}\\ \ell\text{ prime}\end{subarray}}\left(1-\frac{1}{\ell-1}\right).$
Since
$1-\frac{1}{\ell-1}=\left(1-\frac{1}{\ell}\right)\left(1-\frac{1}{(\ell-1)^{2}}\right),$
Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply
$\mathbb{P}(E_{n})\asymp\frac{\log y_{n}}{\log z_{n}}=\frac{1}{a_{n}}\asymp\frac{1}{n}.$
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$. Since the random variables $B_{\ell}$ are mutually independent (Lemma 10) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent. The second Borel–Cantelli lemma therefore gives
$\mathbb{P}(E_{n}\text{ i.o.})=1.$
Let
$R_{n}:=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad F_{n}:=\{R_{n}\leq r_{n}\}.$
Since $w_{\ell}=\frac{1}{\ell}+O(\ell^{-2})$, we have
$\mathbb{E}R_{n}=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{w_{\ell}}{\ell-1}\ll\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell^{2}}\ll\frac{1}{z_{n}}.$
By Markov’s inequality,
$\mathbb{P}(F_{n}^{c})\leq\frac{\mathbb{E}R_{n}}{r_{n}}\ll z_{n}^{-1/2}=r_{n}.$
Thus $\sum_{n}\mathbb{P}(F_{n}^{c})<\infty$, so
$\mathbb{P}(F_{n}^{c}\text{ i.o.})=0.$
For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
$v(\boldsymbol{b}):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
and let
$U_{n}:=\bigcup_{\boldsymbol{b}\in T_{n}}[v(\boldsymbol{b}),\,v(\boldsymbol{b})+r_{n}].$
If $A_{n}\cap E_{n}\cap F_{n}$ occurs, then the block $(y_{n},z_{n}]$ contributes nothing and the tail beyond $z_{n}$ contributes at most $r_{n}$, so $Y\in U_{n}$. Since
$\sum_{n}\mathbb{P}(A_{n}^{c})<\infty,\qquad\sum_{n}\mathbb{P}(F_{n}^{c})<\infty,\qquad\mathbb{P}(E_{n}\text{ i.o.})=1,$
it follows that
$\mu_{f}\!\left(\limsup_{n\to\infty}U_{n}\right)=1.$
On the other hand,
$\sum_{n\geq 1}|T_{n}|\,r_{n}^{s}\ll\sum_{n\geq 1}y_{n}^{Cn}z_{n}^{-s/2}=\sum_{n\geq 1}y_{n}^{Cn-a_{n}s/2}\leq\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Hence
$\mathcal{H}^{s}_{r_{N}}\!\left(\limsup_{n\to\infty}U_{n}\right)\leq\sum_{n\geq N}|T_{n}|\,r_{n}^{s}.$
Letting $N\to\infty$ gives
$\mathcal{H}^{s}\!\left(\limsup_{n\to\infty}U_{n}\right)=0.$
Thus $\mu_{f}$ assigns full mass to a Borel set of zero $s$-dimensional Hausdorff measure, and therefore $\dim_{H}(\mu_{f})\leq s$. Since $s>0$ was arbitrary, $\dim_{H}(\mu_{f})=0$.
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since $T$ is Lipschitz on $\log 2,\infty)$, the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [17, Corollary 2.4]). Therefore $\dim_{H}(\mu_{G})=0$. ∎
The vanishing Hausdorff dimension of the limiting measure explains the clustering visible in Figure 2: the purely singular law forces the values of $c(p)$ to concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
Figure 2: Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Proposition 8 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$. The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$. The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 2.
Theorem 6 (Moment Euler product).
For every real $k>0$,
$\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$ (5)
The product converges absolutely, with the $\ell$-th factor equal to $1+O(k/\ell^{2})$.
Proof.
Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 10), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$. Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
$\mathbb{E}[e^{-kS}]=\prod_{\ell\geq 3}\mathbb{E}[e^{-kZ_{\ell}}]=\prod_{\ell\geq 3}\Bigl(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\Bigr).$
Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (5).
Writing the $\ell$-th factor as $1+a_{\ell}$ with $a_{\ell}=\bigl((\frac{\ell-1}{\ell})^{k}-1\bigr)/(\ell-1)=(e^{-kw_{\ell}}-1)/(\ell-1)$, where $w_{\ell}:=\log(\ell/(\ell-1))=1/\ell+O(1/\ell^{2})$. The bound $|e^{u}-1|\leq|u|\,e^{|u|}$ with $u=-kw_{\ell}$ gives $|a_{\ell}|\leq kw_{\ell}\,e^{kw_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, we obtain $|a_{\ell}|\leq C_{k}/\ell^{2}$ for all $\ell\geq 3$, where $C_{k}>0$ depends only on $k$. Since $\sum_{\ell}\ell^{-2}<\infty$, absolute convergence follows. ∎
Corollary 5 (First and second moments).
  1. (i)
    $\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)}.$
  2. (ii)
    $\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
  3. (iii)
    $\operatorname{Var}(X)=\mathbb{E}[X^{2}]-(\mathbb{E}[X])^{2}$.
Proof.
Immediate from (5) with $k=1$ and $k=2$. ∎
Theorem 7 (Moment determinacy).
The moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ uniquely determines $G$.
Proof.
Since $\operatorname{supp}(G)=[0,\tfrac{1}{2}]$ (Theorem 4), the random variable $X$ is supported on the compact interval $0,\tfrac{1}{2}]$. By the Weierstrass approximation theorem, polynomials are uniformly dense in $C([0,\tfrac{1}{2}])$, so the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines the distribution of $X$ uniquely. ∎
Theorem 8 (Analytic continuation).
The function
$M(s)\;:=\;\mathbb{E}[X^{s}]\;=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
extends to an analytic non-vanishing function on $\{s\in\mathbb{C}:\operatorname{Re}(s)>0\}$, with the Euler product converging uniformly on compact subsets.
Proof.
We first establish the product formula for complex $s$ with $\sigma:=\operatorname{Re}(s)>0$. By Lemma 10, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$. Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$. Since $Z_{\ell}\geq 0$ and $\sigma>0$,
$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq 1$
for every $L$. As $S_{L}\to S:=\sum_{\ell}Z_{\ell}$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$. Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$. Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)$, extending the real-variable identity of Theorem 6 to the half-plane $\operatorname{Re}(s)>0$.
It remains to verify analyticity and non-vanishing. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\{\operatorname{Re}(s)>0\}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [8, Ch. VII]) gives that the product converges uniformly on compact subsets to an analytic function. As $2^{-s}$ is entire, $M(s)$ is analytic on $\{\operatorname{Re}(s)>0\}$.
For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$. To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$. Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{K}/\ell^{2}$). For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{K}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$. The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$. ∎
Proposition 5 (Squared modulus of the characteristic function).
For every $\tau\in\mathbb{R}$,
$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$
Proof.
By Lemma 10, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$. With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
$|\widehat{\nu_{\ell}}(\tau)|^{2}=\frac{(\ell-2)^{2}+2(\ell-2)\cos\alpha+1}{(\ell-1)^{2}}=1-\frac{2(\ell-2)}{(\ell-1)^{2}}(1-\cos\alpha),$
using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. ∎
Proposition 6 (Fourier non-decay implies singularity).
If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular, and consequently $G$ is purely singular.
Proof.
By Lemma 6, $\mu_{f}$ is continuous and of pure type: either absolutely continuous or purely singular. The Riemann–Lebesgue lemma gives: if $\mu_{f}$ is absolutely continuous ($\mu_{f}=h\,\mathrm{d}t$, $h\in L^{1}$), then $\widehat{\mu_{f}}(\tau)=\widehat{h}(\tau)\to 0$. Contrapositively, $\limsup|\widehat{\mu_{f}}(\tau)|>0$ implies $\mu_{f}$ is not absolutely continuous, hence purely singular. For the conclusion about $G$: by Corollary 2, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$, and $G$ is the distribution function of $X$. The map $T(u)=e^{-u}$ is locally Lipschitz on $0,\infty)$, so the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets. Hence pure singularity of $\mu_{f}$ implies pure singularity of $G$ (cf. the proof of Corollary 2). ∎
Figure 3: Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 5 using all primes $\ell\leq 10^{7}$. The top panel uses a linear scale and the bottom panel a log-log scale. The function oscillates but its local averages decrease across successive decades of $\tau$.
Remark 2 (Numerical evidence on Fourier decay).
We evaluated $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 5 with all primes $\ell\leq 10^{7}$. Figure 3 displays the result for $\tau\in[1,10^{6}]$. The function oscillates wildly, with local maxima near $0.14$ persisting at moderately large $\tau$. However, the mean value within each decade decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, $0.03$ for $10^{2},10^{3}]$, $0.02$ for $10^{3},10^{4}]$, $0.01$ for $10^{4},10^{5}]$, and $0.008$ for $10^{5},10^{6}]$. A power-law fit to the envelope of local maxima yields $|\widehat{\mu_{f}}(\tau)|^{2}\approx\tau^{-0.19}$. This is consistent with decay to 0, which would mean $\mu_{f}$ is a Rajchman measure; establishing this rigorously remains open (Section 8).
6.1 Endpoint asymptotics at $1/2$
The value $1/2$ is the supremum of the support of $G$ but is never attained (Remark 1). The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the rejection-sampling overhead $1/c(p)$ is close to its minimum value of $2$. We now determine this rate precisely.
Proposition 7 (Endpoint asymptotic at $1/2$).
Let $G$ be the limiting distribution function from Theorem 4, and let $X$ be a random variable with distribution $G$. Define
$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr)\approx 0.7413.$
Then, as $\varepsilon\downarrow 0$,
$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}.$
Proof.
By Lemma 10,
$X\overset{d}{=}\frac{1}{2}e^{-S},\qquad S:=\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}Z_{\ell},$
where the $Z_{\ell}$ are independent and satisfy
$\mathbb{P}(Z_{\ell}=0)=\frac{\ell-2}{\ell-1},\qquad\mathbb{P}\!\left(Z_{\ell}=\log\frac{\ell}{\ell-1}\right)=\frac{1}{\ell-1}.$
Set
$M(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}.$
For $u>0$ one has
$\frac{u}{1+u}<\log(1+u)\leq u.$
Applying this with $u=1/(\ell-1)$ gives
$\frac{1}{\ell}<\log\frac{\ell}{\ell-1}\leq\frac{1}{\ell-1}.$
Hence, if $\ell\leq(2t)^{-1}$, then
$Z_{\ell}\neq 0\;\Longrightarrow\;Z_{\ell}>\frac{1}{\ell}\geq 2t>t,$
so the event $\{S\leq t\}$ forces $Z_{\ell}=0$ for every odd prime $\ell\leq(2t)^{-1}$. Therefore
$\mathbb{P}(S\leq t)\leq M\!\bigl((2t)^{-1}\bigr).$ (6)
For the lower bound, set
$z_{t}:=\frac{\log(1/t)}{t}\qquad(0<t<e^{-1}).$
By independence,
$\mathbb{P}(S\leq t)\geq M(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$ (7)
Since $Z_{\ell}\leq 1/(\ell-1)$ almost surely,
$\mathbb{E}\!\left[\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\right]=\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\leq\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{(\ell-1)^{2}}\ll\frac{1}{z_{t}}=\frac{t}{\log(1/t)}.$
Thus Markov’s inequality gives
$\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}>t\right)\ll\frac{1}{\log(1/t)}=o(1),$
and (7) becomes
$\mathbb{P}(S\leq t)\geq M(z_{t})\,(1-o(1)).$ (8)
Now
$M(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 1 gives
$\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\sim\frac{2e^{-\gamma}}{\log y}.$
Hence
$M(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6) and (8) imply
$\mathbb{P}(S\leq t)\sim\frac{\kappa}{\log(1/t)}\qquad(t\downarrow 0).$
Since $\mu_{f}$ is continuous by Proposition 3, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(S<-\log(1-2\varepsilon)\right).$
As $\varepsilon\downarrow 0$,
$\log(1-2\varepsilon)=2\varepsilon+O(\varepsilon^{2}),$
so $\log(1/(-\log(1-2\varepsilon)))\sim\log(1/\varepsilon)$. The stated asymptotic follows. ∎
Remark 3 (Twin prime constant).
The constant $\kappa$ equals $C_{2}\,e^{-\gamma}$, where $C_{2}=2\prod_{p\geq 3}(1-(p-1)^{-2})\approx 1.3203$ is the twin prime constant of Hardy and Littlewood [22] (denote $2C_{2}$ therein). This follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $C_{2}/2$.
Corollary 6 (Endpoint modulus at $1/2$).
As $\varepsilon\downarrow 0$,
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=1-G\!\left(\frac{1}{2}-\varepsilon\right)\asymp\frac{1}{\log(1/\varepsilon)}.$
In particular, for every $\alpha>0$, one does not have
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha})\qquad(\varepsilon\downarrow 0),$
so $G$ is not Hölder of any positive order at the endpoint $1/2$.
Proof.
The two-sided order follows immediately from Proposition 7. If for some $\alpha>0$ one had
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha}),$
then Proposition 7 would force
$\frac{1}{\log(1/\varepsilon)}=O(\varepsilon^{\alpha}),$
which is impossible because
$\varepsilon^{\alpha}\log(1/\varepsilon)\to 0\qquad(\varepsilon\downarrow 0).$
7 Explicit Lower Bounds via the Number of Distinct Prime Factors
Theorem 9 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
Let $p\geq 3$ be a prime, and let $\omega=\omega(p-1)$ denote the number of distinct prime factors of p1p-1. Then
$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\qquad P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779,$
where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
Proof.
Write the distinct prime factors of p1p-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$, we have $P(p)\geq P(3)=P_{3}$. Therefore, by (1),
$c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
as claimed. ∎
Corollary 7 (Matching lower bound).
There exists an absolute constant $c>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;\frac{c}{\log\log x}.$
Consequently,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 4: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ third theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 9 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
Corollary 8 (Pointwise universal lower bound).
Let
$m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
Then for every prime $p\geq 3$,
$c(p)\;\geq\;P_{3}\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $p$,
$c(p)\geq\frac{c_{0}}{\log\log p}.$
Proof.
By Theorem 9,
$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Also,
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1,$
so $\omega(p-1)\leq m(p-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
which proves the first claim.
Finally, by Mertens’ third theorem and Lemma 2,
$\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(p-1)}}\asymp\frac{1}{\log\log p},$
as $p\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(p)\ll\log\log p$. ∎
Remark 4 (On effectivity).
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [34, Theorems 4 and 7], Dusart [12, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [41, Theorem 1.1].
Remark 5 (Sharper bounds for specific primes).
Direct computation of the Euler product gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$. For the NIST moduli $p=3329$ and $p=8380417$ [31, 32], $\prod_{j\geq 1}(1-p^{-j})>0.9996$ and >0.9999998>0.9999998, respectively.
Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 9, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
$\omega(p-1)\leq K$ $B(K)$
11 $P_{3}/2\approx 0.280063$
55 $\approx 0.116390$
$10$ $\approx 0.0884704$
$15$ $\approx 0.0776918$
2020 $\approx 0.0715828$
Table 1: Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 9. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
Remark 6 (On the case $\omega(p-1)=1$).
If $\omega(p-1)=1$ and $p$ is an odd prime, then p1p-1 is an even prime power, hence $p-1=2^{m}$ for some $m\geq 1$. Equivalently, $p$ is a Fermat prime. In this case
$\frac{\varphi(p-1)}{p-1}=\frac{1}{2},\qquad c(p)=\frac{1}{2}\,P(p),\qquad P(p):=\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Hence $c(p)\to 1/2$ along any infinite subsequence of Fermat primes, so $B(1)$ is not tight for Fermat primes.
Corollary 9 (Explicit lower bounds for large primes).
Let $p$ be a prime with $p>2^{30}$. Then
$c(p)>\frac{0.56}{\log p}.$
Moreover, if $\omega(p-1)\leq 10$, then
$c(p)>0.088.$
Proof.
The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 9), since
$B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
For the general bound, since every prime divisor of p1p-1 is at most p1p-1,
$\frac{\varphi(p-1)}{p-1}=\prod_{r\mid(p-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq p-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr).$
By Dusart’s explicit product estimate [12, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr).$
Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
Also,
$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>0.9999.$
Therefore
$c(p)=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(p-1)}\;>\;\frac{0.56}{\log p},$
using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, and $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
$\frac{1}{c(p)}\;<\;\frac{\log p}{0.56}\;<\;1.79\log p.$
Remark 7.
By the shifted-prime Erdős–Kac theorem (Barban [3]; see also Halberstam [20] and Elliott [13, Ch. 12]), extending the classical result of Erdős and Kac [15], $(\omega(p-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal over primes $p\leq x$. For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes.
Standard $p$ generic $B(\omega(p-1))$ bound sharper factorization bound
ML-KEM $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
ML-DSA $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
Table 2: The prime moduli appearing in FIPS 203 and FIPS 204 [31, 32]. The third column gives the generic bound $B(\omega(p-1))$ from Theorem 9, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorization of p1p-1 together with $P(p)\geq 1-\frac{1}{p-1}$.
For the NIST-standardized moduli [31, 32], the exact factorizations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$ together with $P(p)\geq 1-1/(p-1)$ give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$. The certified evaluation of Proposition 8 yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorization bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
Context $p$ $\omega(p-1)$ $1/c(p)$
ML-KEM (FIPS 203) $3329$ 22 $2.17$
ML-DSA (FIPS 204) $8380417$ 44 $3.41$
NewHope / FALCON [2, 19] $12289$ 22 $3.00$
BFV/BGV FHE (small) [5, 18, 6] $786433$ 22 $3.00$
BFV/BGV FHE (medium) [5, 18, 6] $2013265921$ 33 $3.75$
NTTRU / BLISS [27, 11] $7681$ 33 $3.75$
64-bit NTT prime [37] $2^{64}-2^{32}+1$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
Table 3: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
Remark 8 (Concrete example).
For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime”; not to be confused with the Ed448-Goldilocks prime [21]), used in lattice cryptography [37], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
Proposition 8 (Certified evaluation of $c(p)$).
Given a prime $p\geq 3$ and the complete factorization $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
  1. (i)
    Compute the totient ratio $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(1-1/r_{i})$, which is exact as a rational number.
  2. (ii)
    Set the truncation depth $J:=\lceil D\log 10/\!\log p\rceil$ and compute the partial Euler product $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, again exact in rational arithmetic.
  3. (iii)
    Return the certified interval
    $\frac{\varphi(p-1)}{p-1}\!\left(P_{J}(p)-\frac{p^{-J}}{p-1}\right)\;\leq\;c(p)\;\leq\;\frac{\varphi(p-1)}{p-1}\,P_{J}(p).$
The width of this certified interval is at most $p^{-J}/(2(p-1))$. Moreover, in step (ii) the exact rational $P_{J}(p)$ has denominator $p^{J(J+1)/2}$, so the intermediate numerators and denominators have $O(J^{2}\log p)$ bits.
Proof.
Steps (i) and (ii) use k1k-1 and J1J-1 multiplications respectively. The certified bound in (iii) follows from the identity $P(p)=P_{J}(p)\prod_{j>J}(1-p^{-j})$ together with $\prod_{j>J}(1-p^{-j})\geq 1-\sum_{j>J}p^{-j}=1-p^{-J}/(p-1)$ and $P(p)\leq P_{J}(p)$, giving $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$. The interval width bound uses $\varphi(p-1)/(p-1)\leq 1/2$. ∎
8 Concluding Remarks
This article studies the PRIM-LWE density constant $c(p)$ unconditionally from extremal, distributional, and computational viewpoints. We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the worst-case rejection-sampling overhead grows only on a doubly logarithmic scale. We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a metrically negligible subset of $0,\tfrac{1}{2}]$ despite having full topological support. Finally, the explicit $\omega(p-1)$-based bounds and the certified evaluation algorithm of Proposition 8 make these asymptotic results effective for concrete cryptographic moduli.
Several questions remain open. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether
$|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity. The numerical evidence of Remark 2 and Figure 3 suggests slow decay, but the oscillatory structure makes rigorous conclusions difficult. Determining a global modulus of continuity of $G$ is also open. Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli; Corollary 8 gives $1/c(p)=O(\log\log p)$ pointwise. Closing the gap between the empirical bound $4$ and the asymptotic order $\log\log p$ remains an interesting open problem.
Declaration of competing interests
The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Research Data Statement
No external datasets were used; all numerical values are reproducible from the explicit formulas and Monte Carlo procedures described in the paper. All empirical values of $c(p)$ in Figures 12 and Table 3 were computed by evaluating the totient ratio $\varphi(p-1)/(p-1)$ from the factorization of p1p-1, followed by the truncated Euler product $P_{J}(p)$ of Proposition 8 with $J=20$; the resulting error is below $10^{-9}$ for every prime $p\geq 3$. The Monte Carlo approximation of $G$ in Figure 1 samples the product distribution of Lemma 10 truncated at the 10,00010{,}000th odd prime $\ell_{N}$. Write $T_{N}:=\sum_{\ell>\ell_{N}}Z_{\ell}$ for the omitted tail and $S_{N}:=\sum_{3\leq\ell\leq\ell_{N}}Z_{\ell}$ for the retained partial sum, so that $X_{N}:=\tfrac{1}{2}e^{-S_{N}}$ approximates $X=\tfrac{1}{2}e^{-S}$ with $S=S_{N}+T_{N}$. The tail mean satisfies $\mathbb{E}[T_{N}]=\sum_{\ell>\ell_{N}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}<10^{-5}$. Since $T_{N}\geq 0$ and $X_{N}\leq\tfrac{1}{2}$, the pointwise coupling bound $0\leq X_{N}-X=X_{N}(1-e^{-T_{N}})\leq T_{N}/2$ holds almost surely, giving $W_{1}(G_{N},G)\leq\tfrac{1}{2}\,\mathbb{E}[T_{N}]<5\times 10^{-6}$ for the Wasserstein-$1$ distance between the truncated and full distribution functions. Markov’s inequality further gives $\mathbb{P}(T_{N}>\varepsilon)<10^{-5}/\varepsilon$ for every $\varepsilon>0$. Figure 3 was produced by evaluating the closed-form Euler product of Proposition 5 over all primes $\ell\leq 10^{7}$. Tables 1 and 2 were computed from the explicit formula $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ of Theorem 9 and the exact factorizations of p1p-1 for the listed moduli, respectively, using only rational arithmetic.