### 06 May 2026 v5

# Primitive-root ratio over prime fields: a shifted-prime distribution of Hausdorff dimension zero and Implications for PRIM-LWE

Abstract.
For a prime $p$, let $c(p)$ denote the limiting fraction of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$. The quantity $c(p)$ is a natural multiplicative deformation of the totient ratio $\varphi(p-1)/(p-1)$ and inherits its distributional behaviour over the primes. Existence and continuity of the limiting law follow from the shifted-prime Erdős–Wintner–Hildebrand framework. We prove the following new results: unconditionally, $\inf_{p}c(p)=0$ and the sharp order is $\min_{p\leq x}c(p)\asymp 1/\log\log x$; the reciprocal satisfies $\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\log\log p)=e^{\gamma}$, and no smaller constant suffices. We give a complete proof, combining an adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy, that the limiting distribution is purely singular, and strengthen this to $\dim_{H}(\mu_{G})=0$, i.e. the limiting measure is carried by a Borel set of Hausdorff dimension zero. The distribution has full topological support $0,\tfrac{1}{2}]$ and admits a Bernoulli-product representation indexed by the odd primes. The moments are given by convergent Euler products, and the Mellin transform $\mathbb{E}[X^{s}]$ extends to an entire function of $s$, non-vanishing on $\operatorname{Re}(s)>0$. Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ with an explicit constant $\kappa$. The quantity $1/c(p)$ equals the dimension-uniform expected rejection-sampling overhead in the reduction from learning with errors (LWE) to PRIM-LWE in lattice-based cryptography. The explicit bounds yield concrete overhead estimates for all primes appearing in current NIST post-quantum standards and representative NTT-friendly moduli.
2020 Mathematics Subject Classification:
11N60, 11A07, 11K65, 28A78, 94A60
1. Introduction
The distribution of the Euler totient ratio $\varphi(n)/n$ over the positive integers was first investigated by Schoenberg [30], who showed that the sequence possesses a continuous asymptotic distribution function. The existence of such a distribution was placed on a rigorous probabilistic footing by the Erdős–Wintner theorem [11], which identifies the class of real additive functions admitting a limiting law. Erdős [10] subsequently proved that this distribution is purely singular with respect to Lebesgue measure, and the Jessen–Wintner zero–one law [18] supplies the underlying dichotomy. The shifted-prime analogue—the behaviour of $\varphi(p-1)/(p-1)$ as $p$ ranges over primes—was addressed by Kátai [20] and, in definitive form, by Hildebrand [16], who extended the Erdős–Wintner framework to additive functions evaluated on shifted primes. Deshouillers and Hassani [7] refined this picture by establishing continuity in arithmetic progressions and proving left non-differentiability of the distribution function at every value $\varphi(m)/m$ with $m$ even.
The quantity
$c(p)\;:=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
arises as the limiting density of $n\times n$ matrices over $\mathbb{F}_{p}$ whose determinant is a primitive root modulo $p$; the precise derivation is recalled in Section 2. Writing $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})$, one has $P(p)\to 1$ as $p\to\infty$, so the behaviour of $c(p)$ is governed by the classical totient ratio $\varphi(p-1)/(p-1)$, and the distributional theory of $c(p)$ reduces to the shifted-prime Erdős–Wintner–Hildebrand framework. In this sense $c(p)$ is a natural multiplicative deformation of $\varphi(p-1)/(p-1)$.
The open question whether $\inf_{p}c(p)=0$ was raised in [31, Remark 3(ii)]; the cryptographic implications of the explicit bounds are developed in Section 8.
Main results
We consolidate our principal findings into four theorems.
[…]
$\min_{p\leq x}c(p)\;\asymp\;\frac{1}{\log\log x}\qquad(x\to\infty).$
The extremal order of the reciprocal is sharp:
$\limsup_{p\to\infty,\,p\textup{ prime}}1/(c(p)\,\log\log p)=e^{\gamma}\approx 1.781.$
In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and no smaller constant suffices. The limsup is realised asymptotically along the least primes $p_{k}^{*}\equiv 1\pmod{N_{k}}$, whose existence is guaranteed by Dirichlet’s theorem and whose size is controlled by Linnik’s theorem.
See Theorem 3.3, Proposition 4.2, Corollary 7.2, and Theorem 7.4. Remark 7.5 discusses effectivity.
Theorem B (Limiting law).The quantity $c(p)$ possesses a continuous limiting distribution $G$ over the primes, coinciding with the distribution of $\varphi(p-1)/(p-1)$. The topological support of $G$ is the full interval $0,\tfrac{1}{2}]$, and $G$ admits a Bernoulli-convolution representation: if $B_{\ell}$ are independent Bernoulli random variables with $\mathbb{P}(B_{\ell}=1)=1/(\ell-1)$ for each odd prime $\ell$, then
$X\;\stackrel{{\scriptstyle d}}{{=}}\;\tfrac{1}{2}\prod_{\ell\geq 3,\;\ell\textup{ prime}}\Bigl(1-\frac{1}{\ell}\Bigr)^{\!B_{\ell}}.$
See Theorems 5.5 and 5.8, Corollary 5.14, and Lemma 6.1. Equivalently, the logarithmic transform of the distribution is an infinite convolution of two-point measures (Lemma 6.1).
Theorem C (Hausdorff dimension zero and singularity).The Hausdorff dimension of the limiting measure $\mu_{G}$ is zero: $\dim_{H}(\mu_{G})=0$. In particular, $\mu_{G}$ is purely singular with respect to Lebesgue measure.
See Theorem 6.2 and Section 5.3.
Theorem D (Mellin transform and endpoint law).For every $s\in\mathbb{C}$, the Mellin transform $M(s)=\mathbb{E}[X^{s}]$ is given by a convergent Euler product; it extends to an entire function of $s$, non-vanishing on $\{\operatorname{Re}(s)>0\}$. Near the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$, where $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ and $\mathfrak{S}_{2}$ is the Hardy–Littlewood twin-prime singular series.
See Theorems 6.3 and 6.4, Proposition 6.9, and Corollary 6.10. The twin-prime singular series enters because the event that no small prime $\ell$ divides p1p-1 forces $\varphi(p-1)/(p-1)$ close to $1/2$. The resulting local factors $1-1/(\ell-1)^{2}$ are precisely those governing the density of twin primes in the Hardy–Littlewood heuristic [15].
As a consequence, explicit lower bounds on $c(p)$ via $\omega(p-1)$ yield effective estimates for any prime whose factorisation of p1p-1 is known; see Section 7.
Davenport [6] independently studied the distribution of $\sigma(n)/n$ over the integers, where $\sigma$ denotes the sum-of-divisors function; for $\varphi(n)/n$ the topological support is the full interval $[0,1]$ and the global modulus of continuity remains unknown (see Section 9). Hildebrand’s proof [16] of the shifted-prime Erdős–Wintner theorem builds on the equidistribution of primes in arithmetic progressions on average established by Bombieri [3]; see also Tenenbaum [32, 33]. Tenenbaum [32] remarked that the pure singularity of the shifted-prime distribution can be obtained by adapting an exercise solution in [34, Exercise 2.4] (corresponding to [33, Exercise 256]), and Deshouillers–Hassani [7] cite the same exercise, but neither source carries out the argument. We give a complete proof in Section 5.3 below, combining an explicit adaptation of Erdős’s argument with the Jessen–Wintner pure-type dichotomy [18]. We strengthen this to the metric result $\dim_{H}(\mu_{G})=0$ in Section 6.1. Wintner’s convergence theory [36]—the founding document for infinite convolutions of discrete measures—supplies the underlying framework. The Hausdorff dimension of Bernoulli convolutions has a substantial literature; see Kahane [19] for an early observation that Erdős-type arguments yield Hausdorff dimension estimates, and Peres, Schlag, and Solomyak [26] and Peres–Solomyak [27] for further dimensional results. The classical Bernoulli convolution involves a fixed contraction ratio. The infinite convolution arising here involves varying ratios $\log(\ell/(\ell-1))$ indexed by the odd primes, so the standard Pisot–Salem machinery does not apply directly. Billingsley [2] and Elliott [13] provide modern probabilistic perspectives.
The extremal behaviour of $\varphi(n)/n$ dates to Mertens [23], whose third theorem gives $\prod_{\ell\leq x}(1-1/\ell)\sim e^{-\gamma}/\log x$; the systematic study of extremal values of arithmetic functions was initiated by Ramanujan [28]. Over the primes, primes $p\equiv 1\pmod{N_{k}}$—for which p1p-1 has many small prime factors—realise the smallest values of the totient ratio, and the size of the least such prime is controlled by Linnik’s theorem [21]. Aivazidis and Sofos [1] considered the density of maximal-order elements in $\mathrm{GL}_{n}(q)$ over prime powers. There is also a conceptual kinship with the Artin primitive-root conjecture (see Hooley [17] for a conditional proof under GRH): while Artin’s conjecture concerns the behaviour of a fixed base as $p$ varies, the present work concerns the aggregate density $c(p)$ in the complementary regime where the base varies uniformly. Both problems ultimately depend on the factorisation of p1p-1.
Organisation
Section 2 recalls the derivation of $c(p)$ and the open question from [31]. Section 3 resolves this question unconditionally, and Section 4 gives quantitative refinements via Linnik’s theorem. Section 5 develops the distributional theory: existence and continuity of the limiting law (Section 5.1), transfer from $\varphi(p-1)/(p-1)$ to $c(p)$ (Section 5.2), and the complete proof of pure singularity (Section 5.3). Section 6 establishes the convolution structure and Hausdorff dimension zero (Section 6.1), the moment Euler product and Mellin transform (Section 6.2), the Fourier structure (Section 6.3), and the endpoint asymptotic (Section 6.4). Section 7 provides explicit lower bounds, Section 8 applies these bounds to the PRIM-LWE reduction in lattice-based cryptography, and Section 9 concludes by discussing two open questions.
2. Background and the Open Question
Throughout, $\log$ denotes the natural logarithm. We write $\omega(n)$ for the number of distinct prime factors of $n$ and $p_{k}$ for the $k$-th prime.
Define (cf. [31])
$M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})\;=\;\bigl\{\,M\in M_{n}(\mathbb{F}_{p}):\det(M)\text{ generates }\mathbb{F}_{p}^{*}\,\bigr\}.$
The determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{p})\longrightarrow\mathbb{F}_{p}^{\times}$
is a surjective group homomorphism, so a fraction $\varphi(p-1)/(p-1)$ of invertible matrices have primitive-root determinant. Since
$|\mathrm{GL}_{n}(\mathbb{F}_{p})|=p^{n^{2}}\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
(see, e.g., Mullen and Panario [25, Theorem 13.3.3]), it follows that
$c_{n}(p)\;:=\;\frac{|M_{n}^{\mathrm{prim}}(\mathbb{F}_{p})|}{|M_{n}(\mathbb{F}_{p})|}\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{n}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
cf. [31, Corollary 4]. Consequently, the limit
$c(p)\;:=\;\lim_{n\to\infty}c_{n}(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
exists and is strictly positive for each fixed $p$. Since $c_{n}(p)$ is decreasing in $n$, $\inf_{n\geq 1}c_{n}(p)=c(p)$.
Writing $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$ for the prime factorisation of p1p-1, one has
(2.1) $c(p)\;=\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{r_{i}}\Bigr)\;\cdot\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr).$
Since $P(p):=\prod_{j=1}^{\infty}(1-p^{-j})\to 1$ as $p\to\infty$, the behaviour of $c(p)$ over the primes is governed by $\varphi(p-1)/(p-1)$, and the distributional theory in Section 5 reduces to the classical shifted-prime distribution of this ratio.
In [31, Remark 3(ii)], Sehrawat, Yeo, and Desmedt observed that if there exist infinitely many primorial primes $p=\prod_{i=1}^{k}p_{i}+1$, then $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(p_{i}-1)/p_{i}\to 0$ by Mertens’ third theorem, which would imply $\inf_{p}c(p)=0$. They left open whether this can be proved unconditionally; see [4] for computational results on primorial primes.
3. Unconditional Proof that $\inf_{p}c(p)=0$
The argument requires only two classical results: Dirichlet’s theorem on primes in arithmetic progressions [8] and the following.
Theorem 3.1 (Mertens’ Third Theorem [23]; see [24, Theorem 2.7(e)], 1874).
As $x\to\infty$,
$\prod_{\begin{subarray}{c}\ell\leq x\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{\ell}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log x},$
where $\gamma\approx 0.5772$ is the Euler–Mascheroni constant.
Lemma 3.2 (Lower bound for the matrix-count Euler product).
For all primes $p\geq 3$,
$\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\geq\;\frac{1}{2}.$
Proof.
By the finite-product inequality $\prod(1-a_{j})\geq 1-\sum a_{j}$ (immediate by induction when $0\leq a_{j}\leq 1$) and the geometric series $\sum_{j\geq 1}p^{-j}=1/(p-1)\leq 1/2$. ∎
Theorem 3.3 (Unconditional vanishing of $\inf c(p)$).
The following identities hold:
$\inf_{p\textup{ prime}}c(p)=0,\qquad\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0.$
Proof.
For each $k\geq 1$ define the $k$th primorial
$N_{k}\;=\;\prod_{i=1}^{k}p_{i}.$
Since $\gcd(1,N_{k})=1$, Dirichlet’s theorem guarantees infinitely many primes in the arithmetic progression $\{1+j\,N_{k}:j\geq 1\}$. Fix any such prime $p\equiv 1\pmod{N_{k}}$, so that $N_{k}\mid(p-1)$.
Because every prime $\ell\leq p_{k}$ divides p1p-1, the multiplicative formula for the totient gives
$\frac{\varphi(p-1)}{p-1}\;=\;\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
By Mertens’ third theorem (Theorem 3.1),
$\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since the infinite product satisfies $\prod_{j=1}^{\infty}(1-p^{-j})\leq 1$ for every prime $p$, we obtain
$c(p)\;=\;\frac{\varphi(p-1)}{p-1}\;\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\xrightarrow{k\to\infty}\;0.$
Since for every $k$ there are infinitely many primes $p\equiv 1\pmod{N_{k}}$, the displayed upper bound holds for arbitrarily large primes $p$. Letting $k\to\infty$ gives
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}c(p)=0,$
and the infimum statement follows immediately. ∎
4. Quantitative Refinements
By Linnik’s theorem [21], there exist absolute constants $C_{0},L>0$ such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$.
Let $N_{k}$ denote the $k$-th primorial, $\vartheta$ the Chebyshev function, and $m(x)$ the primorial-counting function:
$N_{k}:=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p,\qquad m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
Lemma 4.1 (Prime-number-theorem consequences for primorials; cf. [24, Theorem 6.9 and Corollary 2.5]).
As $k\to\infty$,
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}.$
As $x\to\infty$,
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x.$
Proof.
The first two assertions follow from $\vartheta(x)\sim x$. For the last two, $N_{m(x)}\leq x<N_{m(x)}p_{m(x)+1}$ gives $\log x\sim\log N_{m(x)}\sim p_{m(x)}$, since $\log p_{m(x)+1}=o(p_{m(x)})$. ∎
Proposition 4.2 (Upper bound for the minimum density up to $x$).
There exists an absolute constant $C>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;\frac{C}{\log\log x}.$
Proof.
Fix absolute constants $C_{0},L>0$ from Linnik’s theorem such that the least prime in any reduced residue class modulo $d$ is at most $C_{0}d^{L}$. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$. Given $x$, let $k$ be the largest positive integer such that
$C_{0}N_{k}^{\,L}\leq x.$
Then
$p^{*}(k)\leq C_{0}N_{k}^{\,L}\leq x.$
Such a $k$ exists for all sufficiently large $x$, and $k\to\infty$ as $x\to\infty$.
Since $\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}$ by the prime number theorem, the inequality $C_{0}N_{k}^{\,L}\leq x$ gives
$L\,\log N_{k}\leq\log x-\log C_{0},$
and hence $p_{k}\ll\log x$. Therefore $\log p_{k}=O(\log\log x)$.
Conversely, the maximality of $k$ forces $C_{0}N_{k+1}^{\,L}>x$, so
$L\log N_{k+1}>\log x-\log C_{0}.$
Since $\log N_{k+1}=\log N_{k}+\log p_{k+1}$, while
$\log N_{k}=\sum_{i=1}^{k}\log p_{i}\sim p_{k}\quad\text{and}\quad\log p_{k+1}\sim\log p_{k}=o(p_{k})=o(\log N_{k}),$
we have $\log N_{k+1}\sim\log N_{k}$. Hence $\log N_{k}\gg\log x$, so $p_{k}\gg\log x$ and therefore $\log p_{k}=\Omega(\log\log x)$. Combined with the upper bound just proved, this yields
$\log p_{k}\asymp\log\log x.$
By construction the least prime $p^{*}(k)\equiv 1\pmod{N_{k}}$ satisfies $p^{*}(k)\leq x$, and the argument in the proof of Theorem 3.3 gives
$c\bigl(p^{*}(k)\bigr)\;\leq\;\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\sim\;\frac{e^{-\gamma}}{\log p_{k}}\;\asymp\;\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\leq\;c\bigl(p^{*}(k)\bigr)\;\ll\;\frac{1}{\log\log x},$
as claimed. ∎
The matching lower bound $\min_{p\leq x,\,p\text{ prime}}c(p)\gg 1/\!\log\log x$, which completes the proof that the order is sharp, is established in Corollary 7.2; its proof requires only the primorial counting argument and Lemma 3.2. In particular, the lower bound is independent of the distributional machinery in Sections 56: it uses the explicit inequality $c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}(1-1/p_{i})$ of Theorem 7.1 together with the primorial comparison $\omega(p-1)\leq m(p-1)$ and Mertens’ third theorem.
5. The Limiting Distribution of $c(p)$
The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework [11] (Tenenbaum [32], Deshouillers–Hassani [7]; see also [34]). We recall these results in Sections 5.1 and 5.2 for completeness. Pure singularity of the shifted-prime law was anticipated by a remark of Tenenbaum [32], who noted that the integer-case argument of Erdős [10] (with solution in [34, Exercise 2.4]) can be adapted. We carry out this adaptation in Section 5.3, providing a complete proof in the shifted-prime setting. The strictly stronger result $\dim_{H}(\mu_{G})=0$ is established in Section 6.1. Throughout this section we write
(5.1) $f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}\,.$
5.1. The distribution of $\varphi(p-1)/(p-1)$
Lemma 5.1 (Verification of the Erdős–Wintner hypotheses).
The function $f$ defined in (5.1) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
The three convergence conditions of the shifted-prime Erdős–Wintner theorem hold:
$\sum_{|f(\ell)|>1}\frac{1}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,$
as does the continuity criterion $\sum_{f(\ell)\neq 0}1/\ell=\infty$. The same convergence conditions and continuity criterion hold for f-f, and the limiting distribution of $g(p-1)=\varphi(p-1)/(p-1)$ is the push-forward of the limiting distribution of f(p1)f(p-1) under $u\mapsto e^{-u}$.
Proof.
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion is satisfied. Since $\log(\varphi(n)/n)=-f(n)$ is also strongly additive with the same absolute values at primes, all three series conditions and the continuity criterion transfer to f-f. The distributional conclusion for $g(n)=\varphi(n)/n$ then follows by applying the continuous map $u\mapsto e^{-u}$. ∎
Proposition 5.2 (Deshouillers–Hassani; special case of [7, Proposition 2.1 and equation (2.1)]).
For each integer $m\geq 1$, there exists a continuous distribution function $G_{m}\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x;m,1)}\#\Bigl\{p\leq x:p\textup{ prime},\ p\equiv 1\pmod{m},\;\frac{\varphi(p-1)}{p-1}\leq\alpha\Bigr\}\\ =G_{m}(\alpha).$
Moreover, the limiting measure is supported on $0,\varphi(m)/m]$.
Proof.
Apply Deshouillers and Hassani [7, Proposition 2.1] to
$g(n):=\frac{\varphi(n)}{n}.$
Since $\log g(\ell)=-\log(\ell/(\ell-1))=-f(\ell)$, Lemma 5.1 (which verifies the convergence conditions for both $f$ and f-f) shows that the three convergence hypotheses of [7, Proposition 2.1] are satisfied, and there exists a distribution function $G_{m}$ such that the displayed limit holds at every continuity point of $G_{m}$.
By the Jessen–Wintner continuity criterion (applied in [7, Proposition 2.1, equation (2.1)]), $G_{m}$ is continuous if
$\sum_{\begin{subarray}{c}\ell\equiv 1\pmod{m}\\ g(\ell)\neq 1\end{subarray}}\frac{1}{\ell}$
diverges. (The full criterion requires $\sum_{\ell\nmid m,\,g(\ell)\neq 1}1/\ell=\infty$, which holds a fortiori since $g(\ell)\neq 1$ for every prime.) Here $g(\ell)=1-\frac{1}{\ell}\neq 1$ for every prime $\ell$, and
$\sum_{\ell\equiv 1\pmod{m}}\frac{1}{\ell}=\infty$
by the divergence of the prime harmonic series in every reduced residue class (see, e.g., [24, Corollary 4.10]). Thus $G_{m}$ is continuous, so the displayed limit holds for every $\alpha\in\mathbb{R}$.
If $p\equiv 1\pmod{m}$, then every prime divisor of $m$ also divides p1p-1, so
$\frac{\varphi(p-1)}{p-1}=\prod_{\ell\mid(p-1)}\Bigl(1-\frac{1}{\ell}\Bigr)\leq\prod_{\ell\mid m}\Bigl(1-\frac{1}{\ell}\Bigr)=\frac{\varphi(m)}{m},$
which yields the support claim. ∎
We shall also use the following standard Siegel–Walfisz (see [24, Corollary 11.21] or [33, Theorem II.8.17]) estimate in a range of polylogarithmic moduli.
Theorem 5.3 (Siegel–Walfisz).
For every fixed $A>0$ there exists $c_{A}>0$ such that uniformly for integers $m\leq(\log x)^{A}$ and reduced residue classes $a\pmod{m}$,
$\pi(x;m,a)=\frac{\operatorname{Li}(x)}{\varphi(m)}+O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr),$
where
$\pi(x;m,a)=\#\{p\leq x:p\textup{ prime},\ p\equiv a\pmod{m}\}$.
Lemma 5.4 (Density of even totient ratios).
The set
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}$
is dense in $\left[0,\tfrac{1}{2}\right]$.
Proof.
Since $\varphi(m)/m$ depends only on the distinct prime divisors of $m$, we have
$\left\{\frac{\varphi(m)}{m}:m\geq 1\text{ even}\right\}=\left\{\frac{1}{2}\prod_{r\in S}\left(1-\frac{1}{r}\right):S\subset\{\text{odd primes}\},\ S\text{ finite}\right\}.$
Enumerate the odd primes as $r_{1}<r_{2}<\cdots$, and set
$a_{j}:=\log\frac{r_{j}}{r_{j}-1}.$
Then $a_{j}>0$, $a_{j}\downarrow 0$, and
$a_{j}=\frac{1}{r_{j}}+O\!\left(\frac{1}{r_{j}^{2}}\right).$
Since $\sum_{j}1/r_{j}$ over the odd primes diverges, it follows that
$\sum_{j=1}^{\infty}a_{j}=+\infty.$
We claim that the set of finite subsums
$\left\{\sum_{j\in J}a_{j}:J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $0,\infty)$. Indeed, fix $y\geq 0$ and $\varepsilon>0$. Choose $J_{0}$ so large that $a_{j}<\varepsilon$ for all $j\geq J_{0}$. Since $\sum_{j\geq J_{0}}a_{j}=\infty$, there exists $N\geq J_{0}$ minimal such that
$\sum_{j=J_{0}}^{N}a_{j}>y.$
By minimality,
$\sum_{j=J_{0}}^{N-1}a_{j}\leq y<\sum_{j=J_{0}}^{N}a_{j}=\sum_{j=J_{0}}^{N-1}a_{j}+a_{N}.$
Because $a_{N}<\varepsilon$, it follows that
$y-\varepsilon<\sum_{j=J_{0}}^{N-1}a_{j}\leq y.$
Thus a finite subsum lies within $\varepsilon$ of $y$, proving the claim.
Exponentiating the negative of these finite subsums therefore shows that
$\left\{\frac{1}{2}\exp\!\left(-\sum_{j\in J}a_{j}\right):J\subset\mathbb{N}\ \text{finite}\right\}=\left\{\frac{1}{2}\prod_{j\in J}\left(1-\frac{1}{r_{j}}\right):J\subset\mathbb{N}\ \text{finite}\right\}$
is dense in $(0,1/2]$. Since the finite subsums can be arbitrarily large, 0 is also a limit point. Hence the displayed set is dense in $[0,1/2]$, as claimed. ∎
Theorem 5.5 (Limiting distribution and full support over primes).
There exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}\;=\;G(\alpha).$
The limiting law has support exactly $0,\tfrac{1}{2}]$, with $G(\alpha)=0$ for $\alpha\leq 0$ and $G(\alpha)=1$ for $\alpha\geq 1/2$. In particular, $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Apply Proposition 5.2 with $m=2$. Since every odd prime satisfies $p\equiv 1\pmod{2}$, we have
$\pi(x;2,1)=\pi(x)-1\sim\pi(x),$
so $G_{2}$ is the unconditional limiting distribution of $\varphi(p-1)/(p-1)$ over all primes. Writing $G:=G_{2}$ (denoted $F$ in [7, Theorem 1.1]), there exists a continuous distribution function $G\colon\mathbb{R}\to[0,1]$ such that for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\text{ prime},\;\varphi(p-1)/(p-1)\leq\alpha\bigr\}=G(\alpha).$
For every odd prime $p$, the integer p1p-1 is even, so
$0<\frac{\varphi(p-1)}{p-1}\leq\frac{1}{2}.$
The exceptional prime $p=2$ is negligible in the limit. Hence the limiting law is carried by $0,\tfrac{1}{2}]$, and therefore
$G(\alpha)=0\qquad(\alpha\leq 0),\qquad G(\alpha)=1\qquad(\alpha\geq 1/2).$
Let $\mu_{G}$ denote the limiting measure with distribution function $G$. Fix an even integer $m\geq 2$, and write
$x_{m}:=\frac{\varphi(m)}{m}.$
By Deshouillers and Hassani [7, Theorem 1.1], the distribution function $G$ is not left-differentiable at $x_{m}$. Deshouillers and Hassani denote this distribution function by $F$; we write it as $G$. If $x_{m}\notin\operatorname{supp}(\mu_{G})$, then some open interval $I$ containing $x_{m}$ has $\mu_{G}(I)=0$. Hence $G$ is locally constant on $I$, and in particular left-differentiable at $x_{m}$, a contradiction. Therefore $x_{m}\in\operatorname{supp}(\mu_{G})$ for every even $m$.
By Lemma 5.4, the support contains a dense subset of $0,\tfrac{1}{2}]$. Since the support is closed and already lies inside $0,\tfrac{1}{2}]$, it follows that
$\operatorname{supp}(\mu_{G})=[0,\tfrac{1}{2}].$
Now let $0\leq a<b\leq 1/2$. Choose any $x\in(a,b)$. Since $x\in\operatorname{supp}(\mu_{G})$, every neighbourhood of $x$ has positive $\mu_{G}$-measure. Taking $r>0$ so small that $(x-r,x+r)\subset(a,b)$, we obtain
$\mu_{G}((a,b])\geq\mu_{G}((x-r,x+r))>0.$
Hence
$G(b)-G(a)=\mu_{G}((a,b])>0,$
so $G$ is strictly increasing on $0,\tfrac{1}{2}]$. The displayed bounds $0<G(\tau)<1$ for $0<\tau<1/2$ follow immediately from this and the endpoint values $G(0)=0$, $G(1/2)=1$. ∎
Remark 5.6 (Alternative support argument).
Full support also follows from the Bernoulli-convolution representation established in Section 6.1: the support of the infinite convolution $\mu_{f}$ of Lemma 6.1 is the closure of the set of partial sums $\log 2+\sum_{\ell\in S}\log(\ell/(\ell-1))$ over finite sets $S$ of odd primes, and the density argument of Lemma 5.4 shows this closure equals $\log 2,\infty)$. The proof above uses instead the non-differentiability result of Deshouillers–Hassani, which connects the support question to the local analytic behaviour of $G$.
For the classical distribution of $\varphi(n)/n$ over the integers, the topological support is $[0,1]$: the argument of Lemma 5.4, applied over all primes and without the forced factor $1/2$, gives density of $\{\varphi(n)/n:n\geq 1\}$ in $[0,1]$, and the Erdős–Wintner theorem [11] (see also Schoenberg [30]) identifies this closure with the support. The global modulus of continuity remains open; see Section 9.
5.2. From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 5.7 (Perturbation lemma).
Let $X(p)\geq 0$ and $a(p)>0$ be functions of the primes. Assume that $X(p)$ has a continuous limiting distribution function $G$ over the primes, i.e.
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\ X(p)\leq\alpha\bigr\}=G(\alpha)\qquad(\alpha\in\mathbb{R}).$
If $a(p)\to 1$ as $p\to\infty$ through the primes, then $a(p)X(p)$ has the same limiting distribution $G$.
Proof.
For $\alpha<0$ both sides are zero since $X(p)\geq 0$ and $a(p)>0$, so the claim is immediate. We may therefore assume $\alpha\geq 0$.
Fix $0<\varepsilon<1$. For all sufficiently large primes, $1-\varepsilon<a(p)<1+\varepsilon$, so
$\{X(p)\leq\alpha/(1+\varepsilon)\}\;\subseteq\;\{a(p)X(p)\leq\alpha\}\;\subseteq\;\{X(p)\leq\alpha/(1-\varepsilon)\}$
up to finitely many exceptions. Taking densities as $x\to\infty$ and letting $\varepsilon\downarrow 0$, continuity of $G$ gives the claim. ∎
Theorem 5.8 (Transfer to the primitive-root determinant density).
The quantity $c(p)$ has the same limiting distribution as $\varphi(p-1)/(p-1)$. That is, for every $\alpha\in\mathbb{R}$,
$\lim_{x\to\infty}\;\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;c(p)\leq\alpha\bigr\}\;=\;G(\alpha),$
where $G$ is the distribution function from Theorem 5.5.
Proof.
Write
$X(p):=\frac{\varphi(p-1)}{p-1},\qquad P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr),$
so that $c(p)=X(p)P(p)$.
For every prime $p\geq 2$,
[…]
hence
$P(p)=1-O\!\left(\frac{1}{p}\right)\qquad(p\to\infty\text{ through the primes}).$
By Theorem 5.5, $X(p)$ has continuous limiting distribution function $G$. Therefore Lemma 5.7 applies with $a(p)=P(p)$, and shows that $c(p)=a(p)X(p)$ has the same limiting distribution $G$. ∎
5.3. Pure singularity
Having established that $c(p)$ shares the limiting distribution of $\varphi(p-1)/(p-1)$, we now investigate the fine structure of this law.
Note. The strictly stronger result $\dim_{H}(\mu_{G})=0$—that is, the limiting measure is carried by a Borel set of Hausdorff dimension zero—is established independently in Section 6.1 by a different method (a Borel–Cantelli covering argument exploiting the independence structure of the Bernoulli convolution). That result implies pure singularity as an immediate corollary. We include the Erdős-style argument for its independent methodological interest: it provides the first complete adaptation of Erdős’s classical technique [10] to the shifted-prime setting.
The shifted-prime Erdős–Wintner theorem (Theorem 5.9) gives the existence and characteristic function of the limit law for any additive function on $\{p-1\}$ satisfying three convergence conditions; for strongly additive functions it also yields that the law is of pure type and identifies a continuity criterion. Since Lemma 5.1 verifies the hypotheses for $f=\log(n/\varphi(n))$ and the continuity criterion holds, the remaining question is whether the law is absolutely continuous or purely singular. The latter turns out to hold: we prove pure singularity by adapting a classical argument of Erdős to the shifted-prime setting.
Theorem 5.9 (Shifted-prime Erdős–Wintner theorem; Hildebrand [16], Jessen–Wintner [18]).
Let $f$ be a real-valued additive arithmetical function. The following are equivalent:
  1. (i)
    The distribution functions
    $\frac{1}{\pi(x)}\,\#\bigl\{p\leq x:p\textup{ prime},\;f(p-1)\leq z\bigr\}$
    converge weakly, as $x\to\infty$, to a distribution function $G(z)$.
  2. (ii)
    The three series
    (5.2)  $\sum_{\begin{subarray}{c}\ell\textup{ prime}\\ |f(\ell)|>1\end{subarray}}\frac{1}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)^{2}}{\ell},\qquad\sum_{\ell\textup{ prime}}\frac{f^{*}(\ell)}{\ell}$
    all converge, where $f^{*}(n):=f(n)\mathbf{1}_{|f(n)|\leq 1}$.
When these conditions hold, the characteristic function of the limit law is
(5.3) $\widehat{G}(\tau)\;=\;\prod_{\ell\textup{ prime}}\biggl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\biggr).$
Furthermore, if $f$ is strongly additive (i.e. $f(\ell^{\nu})=f(\ell)$ for every prime $\ell$ and every $\nu\geq 1$):
  1. (a)
    (Pure type.) The limit law is of pure type: purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure.
  2. (b)
    (Continuity criterion.) The limit law is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
Proof.
[…]
Concretely, Hildebrand’s Theorem 5 [16] is stated for the distribution functions $\nu_{x}(p;\,f(p+1)\leq z):=(1/\pi(x))\#\{p\leq x:f(p+1)\leq z\}$. Replacing $p+1$ by p1p-1 throughout changes only the residue class in the Bombieri–Vinogradov input:
$\pi(x;q,-1)\;\longrightarrow\;\pi(x;q,1),$
for which the same level of distribution holds (see, e.g., [16]). The resulting theorem for $\{p-1\}$ is therefore identical in statement and proof, and no new estimates are required.
The explicit characteristic function (5.3) is recorded by Tenenbaum [32, formula (3)] for a general nonzero shift $a$, confirming the Euler-product form directly. For a=1a=-1 no prime divides $a$, so the product in (5.3) runs over all primes $\ell$ without restriction.
Pure type and continuity (strongly additive case). Assume now that $f$ is strongly additive; then $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, so each local factor in (5.3) reduces to the characteristic function of an atomic probability measure supported on $\{0,f(\ell)\}$. (For $\ell=2$ the weight at 0 vanishes and the factor degenerates to the point mass $\delta_{f(2)}$; for odd $\ell$ it is a genuine two-point measure.) By Lévy’s uniqueness theorem (see, e.g., [33, Section III.2]), the limit law equals the infinite convolution of these local factors. The Jessen–Wintner theorem [18] (see also [33, Theorem III.2.7(b)]) states that every infinite convolution of discrete probability measures is of pure type, giving (a).
For the continuity criterion (b), since $f$ is strongly additive each local factor reduces to $\frac{\ell-2}{\ell-1}\,\delta_{0}+\frac{1}{\ell-1}\,\delta_{f(\ell)}$, a two-point measure. Let $\sigma_{\ell}$ denote its maximal atom. The Jessen–Wintner theorem gives: the convolution is continuous if and only if $\sum_{\ell}(1-\sigma_{\ell})=\infty$. When $f(\ell)\neq 0$, the atoms at 0 and at $f(\ell)$ are distinct, so $\sigma_{\ell}=\max\!\bigl(\frac{\ell-2}{\ell-1},\,\frac{1}{\ell-1}\bigr)=\frac{\ell-2}{\ell-1}$ for $\ell\geq 3$, whence $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. When $f(\ell)=0$, the two-point measure collapses to $\delta_{0}$ and $1-\sigma_{\ell}=0$. Therefore $\sum_{\ell}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$. ∎
For $y\geq 3$ and each prime $p$, define
$a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y\\ \ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell,$
and write $P^{+}(n)$ for the largest prime factor of $n$.
Lemma 5.10 (Prime density for a prescribed smooth kernel).
Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. The set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
$\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
among the primes, in the sense that
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
Moreover, writing
$Q(y):=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\frac{\ell-2}{\ell-1}\;\asymp\;\frac{1}{\log y},$
we have $\delta_{y}(m)\geq Q(y)/\varphi(m)$.
Proof.
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5.3) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $Q(y)\asymp 1/\!\log y$, write
$\log Q(y)=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem [24, Theorem 2.7(d)] the sum over primes equals $\log\log y+O(1)$, so $Q(y)=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 5.11 (Tail-mean estimate for f(p1)f(p-1)).
Let $f$ be as in (5.1), and for each prime $p$ and parameter $y\geq 3$ define
$T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then, for each fixed $y\geq 3$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$.
Proof.
Since $f(\ell)=1/\ell+O(1/\ell^{2})$, we have $\eta(y)=\sum_{\ell>y}1/(\ell(\ell-1))+O(\sum_{\ell>y}1/\ell^{3})\ll 1/y$. Interchanging sums gives
$\sum_{p\leq x}T_{y}(p)\;=\;\sum_{\begin{subarray}{c}y<\ell<x\\ \ell\text{ prime}\end{subarray}}f(\ell)\,\pi(x;\ell,1).$
Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5.3),
$\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
Since $f(\ell)\leq 1$ for every prime $\ell$ and $\{\text{primes }\ell\in(y,L]\}\leq L$, the total error is
$\sum_{y<\ell\leq L}f(\ell)\,|E(x;\ell,1)|\;\leq\;L\cdot O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr)\;=\;o\bigl(\pi(x)\bigr).$
The main term from this range is $\operatorname{Li}(x)\sum_{y<\ell\leq L}f(\ell)/(\ell-1)=\operatorname{Li}(x)\bigl(\eta(y)+O(1/L)\bigr)$, since the tail $\sum_{\ell>L}f(\ell)/(\ell-1)\ll 1/L$.
Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}$
$\;\ll\;\frac{x}{L}\;+\;\log\log x.$
Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
Lemma 5.12 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let $f$ be the strongly additive function defined in (5.1). The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
Proof.
We adapt the argument of Erdős [10, §3] (see also [34, Exercise 2.4]) to the shifted-prime setting, replacing the integer equidistribution input with the Siegel–Walfisz theorem (Theorem 5.3). Throughout, $y$ is fixed while $x\to\infty$; the moduli arising from the construction divide $M_{y}:=\prod_{\ell\leq y}\ell$ and hence lie in the Siegel–Walfisz range. Only after taking $x\to\infty$ do we let $y\to\infty$.
By Lemma 5.1 and Theorem 5.9, $\mu_{f}$ exists, is continuous, and of pure type. It remains to rule out absolute continuity.
Step 1: Tail estimate. Fix $\delta\in(0,\tfrac{1}{2})$ and set $R:=\lfloor y^{\delta}\rfloor$. Decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$. By Lemma 5.11 and Markov’s inequality,
(5.4) $\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1}\;\xrightarrow{y\to\infty}\;0.$
Step 2: Small-measure set with positive prime density. Let $m_{1}<\cdots<m_{R}$ be the smallest $R$ even squarefree positive integers. Since even squarefree integers have positive asymptotic density (equal to $2/\pi^{2}$), there is an absolute constant $C>0$ with $m_{j}\leq Cj$; in particular $m_{j}\leq CR=C\lfloor y^{\delta}\rfloor\ll y^{1/2}$ since $\delta<\tfrac{1}{2}$, so $P^{+}(m_{j})\leq m_{j}<y$ for all sufficiently large $y$. Because $\varphi(m_{j})\leq m_{j}\leq Cj$, we have $\sum_{j\leq R}1/\varphi(m_{j})\geq C^{-1}\sum_{j\leq R}1/j\gg\log R$. By Lemma 5.10, $\delta_{y}(m_{j})\geq Q(y)/\varphi(m_{j})$ with $Q(y)\asymp 1/\!\log y$, so
(5.5) $\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;Q(y)\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\log R}{\log y}\;=\;\delta+o(1)\;\gg_{\delta}\;1$
for all sufficiently large $y$, where the implied constants are absolute (from Mertens’ theorem and the density of even squarefree integers) and $\log R/\!\log y=\delta+o(1)$ because $R=\lfloor y^{\delta}\rfloor$. Define the closed set $\bar{I}_{y}:=\bigcup_{j=1}^{R}[f(m_{j}),\,f(m_{j})+1/R^{2}]$. Then $\lambda(\bar{I}_{y})\leq 1/R\to 0$ as $y\to\infty$, while for any $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$ one has $f(p-1)\in\bar{I}_{y}$.
Step 3: Contradiction. Since $a_{y}(p-1)$ takes a unique value for each prime $p$, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint. Finite additivity of the natural density therefore gives $\lim_{x\to\infty}\frac{1}{\pi(x)}\#\{p\leq x:a_{y}(p-1)\in\{m_{1},\dots,m_{R}\}\}=\sum_{j\leq R}\delta_{y}(m_{j})$. Removing the primes for which $T_{y}(p)>1/R^{2}$ loses at most the quantity in (5.4), so combining with (5.5): there exists $c_{0}>0$ independent of $y$ such that for all large $y$,
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}\;>\;0,$
the limit being justified by the Portmanteau theorem (see, e.g., [2, Theorem 2.1]): $\bar{I}_{y}$ is a $\mu_{f}$-continuity set because its boundary consists of at most $2R$ points and $\mu_{f}$ is atom-free (Theorem 5.9(b)), so weak convergence of the empirical measures to $\mu_{f}$ gives $\mu_{x}(\bar{I}_{y})\to\mu_{f}(\bar{I}_{y})$. Since $\lambda(\bar{I}_{y})\to 0$ while $\mu_{f}(\bar{I}_{y})\geq c_{0}$, the measure $\mu_{f}$ cannot be absolutely continuous. Being continuous and of pure type, it is purely singular. ∎
Proposition 5.13 (Shifted-prime limiting law for $f$).
Let $f$ be as in (5.1). Then the empirical measures
$\mu_{x}:=\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}\delta_{f(p-1)}$
converge weakly, as $x\to\infty$, to a Borel probability measure $\mu_{f}$ on $0,\infty)$ that is continuous and purely singular with respect to Lebesgue measure.
Proof.
By Lemma 5.1, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Theorem 5.9) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Theorem 5.9.
It remains to rule out absolute continuity. By Lemma 5.12, $\mu_{f}$ is purely singular. ∎
Corollary 5.14 (Structure of the limiting law for $c(p)$).
The ratio $c(p)$ has a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$. Its distribution function $G$ is strictly increasing on $0,\tfrac{1}{2}]$.
Proof.
Combine Theorems 5.5 and 5.8: $c(p)$ has a continuous limiting law with support $0,\tfrac{1}{2}]$ whose distribution function is strictly increasing. It remains to justify pure singularity.
By Proposition 5.13, the shifted-prime law $\mu_{f}$ of f(p1)f(p-1) (with $f$ as in (5.1)) is continuous and purely singular. Since $\varphi(p-1)/(p-1)=e^{-f(p-1)}$ and $T(u):=e^{-u}$ is Lipschitz on $0,\infty)$ (since $|T^{\prime}(u)|=e^{-u}\leq 1$), the push-forward $T_{\#}\mu_{f}$ preserves Lebesgue-null sets. Moreover, $T$ is injective on $\log 2,\infty)$ and $\mu_{f}$ concentrates on this half-line (since $Y\geq\log 2$ a.s.), so the carrying null set of $\mu_{f}$ maps bijectively to a null set carrying $T_{\#}\mu_{f}$. Hence $T_{\#}\mu_{f}$ is purely singular. Since Theorem 5.8 shows that $c(p)$ has the same limiting distribution $G$ as $\varphi(p-1)/(p-1)$, the limiting measure for $c(p)$ is exactly the same measure $\mu_{G}$, and pure singularity carries over. ∎
The strictly stronger metric result $\dim_{H}(\mu_{G})=0$ is established in Theorem 6.2 by an independent method (a Borel–Cantelli covering argument exploiting the Bernoulli-convolution independence structure).
Figure 1 compares the empirical distribution of $c(p)$ for primes $p\leq 10^{6}$ with the limiting law $G$ of Theorem 5.8; the close agreement already visible at this scale provides a useful numerical check on the distributional theory.
Remark 5.15 (Truncation error for numerical approximation).
Truncate the Bernoulli series of Lemma 6.1 at the first $N$ odd primes ($\ell\leq\ell_{N}$) and let $G_{N}$ denote the distribution function of the truncated law $X_{N}:=\tfrac{1}{2}\prod_{3\leq\ell\leq\ell_{N}}(1-1/\ell)^{B_{\ell}}$. Since $|X-X_{N}|\leq\tfrac{1}{2}T_{N}$ almost surely, where $T_{N}:=\sum_{\ell>\ell_{N},\,\ell\textup{ prime}}B_{\ell}\log(\ell/(\ell-1))$, the coupling characterisation of the Wasserstein-1 distance gives
$W_{1}(G_{N},G)\;\leq\;\tfrac{1}{2}\,\mathbb{E}[T_{N}]\;=\;\tfrac{1}{2}\sum_{\begin{subarray}{c}\ell>\ell_{N}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\,.$
For N=10, [^0] N=10{,}000 (so $\ell_{N}=104{,}743$), $\mathbb{E}[T_{N}]<10^{-5}$ and therefore $W_{1}(G_{N},G)<5\times 10^{-6}$. Since $G_{N}$ has an atom of mass $Q(\ell_{N})\approx 0.064$ at $\tfrac{1}{2}$—where $Q(y)$ is the product defined in Lemma 5.10, equal to the probability that every Bernoulli indicator up to $\ell_{N}$ equals zero—while $G$ is continuous, the truncation error is controlled in $W_{1}$, not in the Kolmogorov metric. The plotted theoretical curve is a Monte Carlo approximation to $G_{N}$; the DKW inequality controls the sampling error (see Figure 1).
Figure 1. Empirical CDF of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, overlaid with the theoretical CDF $G$ of Theorem 5.8, approximated by $10^{7}$ Monte Carlo samples of the truncated convolution $G_{N}$ with N=10, [^0] N=10{,}000 odd primes. See Remark 5.15 for truncation and sampling error bounds. Empirical $c(p)$ values use Remark 7.9 with $J=20$.
Remark 5.16 (The value $1/2$: boundary but not attained).
The value $1/2$ is not attained: $c(p)<1/2$ for every prime $p$. Indeed, for $p=2$ one has $\varphi(p-1)/(p-1)=1$, so $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888<1/2$. For odd $p$, one has $\varphi(p-1)/(p-1)\leq 1/2$ together with $\prod_{j\geq 1}(1-p^{-j})<1$. Thus $1/2$ is only a boundary point of the support, not an attained value. Since the support of $G$ is exactly $0,\tfrac{1}{2}]$ (Corollary 5.14), it follows that $\limsup_{p\to\infty,\,p\textup{ prime}}c(p)=1/2$; for every $\varepsilon>0$, the set of primes with $c(p)>1/2-\varepsilon$ has relative density $1-G(1/2-\varepsilon)>0$ among the primes.
Remark 5.17 (Fixed-dimension analogue).
For each fixed $n\geq 1$, $c_{n}(p)=\frac{\varphi(p-1)}{p-1}P_{n}(p)$ with $P_{n}(p):=\prod_{j=1}^{n}(1-p^{-j})\to 1$, so Lemma 5.7 gives $c_{n}(p)$ the same limiting distribution as $\varphi(p-1)/(p-1)$, and the primorial/Mertens argument yields $\min_{p\leq x,\,p\textup{ prime}}c_{n}(p)\asymp 1/\!\log\log x$.
6. Hausdorff Dimension, Moments, and Fourier Structure
The purely singular limiting law established in Section 5 raises a natural question: how thin is the set carrying full mass of the limiting measure? Theorem 6.2 below provides the strictly stronger result $\dim_{H}(\mu_{f})=0$, giving in particular an independent proof of pure singularity. The Erdős-style singularity argument in Section 5.3 is included for its independent methodological interest. Because $\mu_{f}$ decomposes as an infinite convolution of discrete measures (one for each odd prime), one can exploit the resulting independence structure to build efficient coverings and bound the Hausdorff dimension. We show that $\dim_{H}(\mu_{f})=0$, meaning that the mass concentrates on a Borel set of Hausdorff dimension zero even though the topological support of the push-forward $\mu_{G}$ is the full interval $0,\tfrac{1}{2}]$. We then derive an explicit Euler product for every moment of $G$, prove that these moments determine the distribution uniquely, and establish the analytic continuation of the associated Mellin transform. The section concludes with a precise asymptotic for the behaviour of $G$ near the endpoint $1/2$.
Throughout, $f$ is the strongly additive function defined in (5.1), $\mu_{f}$ is the shifted-prime limiting measure of f(p1)f(p-1), and $X$ is a random variable with distribution $G$. By Corollary 5.14, $X\overset{d}{=}e^{-Y}$ where $Y\sim\mu_{f}$. The Bernoulli indicators $B_{\ell}$ and the scaled jumps $Z_{\ell}:=B_{\ell}\,w_{\ell}$ (with $w_{\ell}=\log(\ell/(\ell-1))$) are used interchangeably below; the former appears in the covering construction (Theorem 6.2), the latter in the convolution decomposition (Lemma 6.1).
6.1. Convolution structure and Hausdorff dimension
Lemma 6.1 (Factor structure of $\mu_{f}$).
The measure $\mu_{f}$ is the infinite convolution
$\mu_{f}\;=\;\delta_{\log 2}\;*\;\mathop{{\ast}}_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\nu_{\ell},$
where $\delta_{\log 2}$ is the point mass at $\log 2$ and, for each odd prime $\ell$,
$\nu_{\ell}\;:=\;\frac{\ell-2}{\ell-1}\,\delta_{0}\;+\;\frac{1}{\ell-1}\,\delta_{f(\ell)},\qquad f(\ell)=\log\frac{\ell}{\ell-1}.$
Equivalently, $Y\overset{d}{=}\log 2+\sum_{\ell\geq 3,\;\ell\textup{ prime}}Z_{\ell}$ with the $Z_{\ell}$ independent, $\mathbb{P}(Z_{\ell}=0)=(\ell-2)/(\ell-1)$, $\mathbb{P}(Z_{\ell}=f(\ell))=1/(\ell-1)$.
Proof.
By Theorem 5.9, the characteristic function of $\mu_{f}$ is given by (5.3):
$\widehat{\mu_{f}}(\tau)=\prod_{\ell\text{ prime}}\Bigl(1-\frac{1}{\ell-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(\ell^{\nu})}}{\ell^{\nu}}\Bigr).$
Since $f$ is strongly additive, $f(\ell^{\nu})=f(\ell)$ for all $\nu\geq 1$, whence $\sum_{\nu\geq 1}e^{i\tau f(\ell^{\nu})}\ell^{-\nu}=e^{i\tau f(\ell)}/(\ell-1)$. At each odd prime $\ell$ the local factor equals $\frac{\ell-2}{\ell-1}+\frac{e^{i\tau f(\ell)}}{\ell-1}=\widehat{\nu_{\ell}}(\tau)$; at $\ell=2$ it equals $e^{i\tau\log 2}=\widehat{\delta_{\log 2}}(\tau)$. Since the characteristic function determines a probability measure uniquely, the convolution decomposition follows. ∎
For a Borel probability measure $\mu$ on $\mathbb{R}$, recall the Hausdorff dimension of $\mu$ (see, e.g., Falconer [14, Chapter 2]):
$\dim_{H}(\mu):=\inf\{\dim_{H}(E):E\subset\mathbb{R}\text{ is Borel and }\mu(E)=1\}.$
Theorem 6.2 (Hausdorff dimension zero of the limiting law).
$\dim_{H}(\mu_{f})=0$. Consequently, if $\mu_{G}$ denotes the limiting measure on $0,\tfrac{1}{2}]$ with distribution function $G$, then $\dim_{H}(\mu_{G})=0$.
Proof.
By Lemma 6.1,
$Y\overset{d}{=}\log 2+\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad w_{\ell}:=\log\frac{\ell}{\ell-1},$
where the $B_{\ell}$ are independent Bernoulli random variables with
$\mathbb{P}(B_{\ell}=1)=\frac{1}{\ell-1},\qquad\mathbb{P}(B_{\ell}=0)=\frac{\ell-2}{\ell-1}.$
Fix $s>0$. We shall show that $\mu_{f}$ gives full mass to a Borel set of vanishing $s$-dimensional Hausdorff measure. The argument proceeds in five steps: (1) construct recursively a sequence of super-exponentially growing covering scales; (2) bound the number of prefix configurations via Chebyshev’s inequality; (3) show that empty-block events occur infinitely often by the second Borel–Cantelli lemma; (4) control the tail contribution by Markov’s inequality; (5) assemble the covering and estimate its Hausdorff measure.
Step 1: Recursive covering scales. For $y\geq 3$, let
$I(y):=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right).$
If $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y}$ is a prefix configuration, then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y}=\boldsymbol{b}\bigr)=e^{-I_{y}(\boldsymbol{b})},$
where $I_{y}(\boldsymbol{b})$ denotes the value of $I(y)$ at that configuration. Moreover,
$\mathbb{E}I(y)\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{\log\ell}{\ell}\ll\log y$
and
$\operatorname{Var}(I(y))$ $\leq\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\mathbb{E}\!\left[\left(B_{\ell}\log(\ell-1)+(1-B_{\ell})\log\frac{\ell-1}{\ell-2}\right)^{\!2}\right]$
$\ll\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{(\log\ell)^{2}}{\ell}\;\ll\;(\log y)^{2},$
where the final estimate follows by partial summation and the prime number theorem. Hence there exists an absolute constant $C_{0}>0$ such that
$\mathbb{E}I(y)\leq C_{0}\log y,\qquad\operatorname{Var}(I(y))\leq C_{0}(\log y)^{2}\qquad(y\geq 3).$
Choose $C:=4C_{0}$ and define
$a_{n}:=\left\lceil\frac{4Cn}{s}\right\rceil\qquad(\text{recall }s>0\text{ is fixed}),\qquad z_{n}:=y_{n}^{a_{n}},\qquad r_{n}:=z_{n}^{-1/2}.$
We construct the sequence $(y_{n})$ recursively. Set $y_{1}:=3$. Given $y_{n}$, define $a_{n}$, $z_{n}$, $r_{n}$ as above and put $y_{n+1}:=z_{n}+1$. Then the intervals $(y_{n},z_{n}]$ are pairwise disjoint by construction, $z_{n}=y_{n}^{a_{n}}$ grows super-exponentially, and the summability conditions $\sum r_{n}<\infty$ and $\sum y_{n}^{-Cn}<\infty$ are immediate since $r_{n}=z_{n}^{-1/2}$ and $y_{n}^{-Cn}\leq y_{n}^{-C}$ with $y_{n}\to\infty$ super-exponentially.
Since $a_{n}\geq 4Cn/s$, it follows that
[…]
this inequality drives the convergence of the covering estimate below.
Step 2: Concentration and cardinality of prefix configurations. For each $n$, set $I_{n}:=I(y_{n})$ and
$A_{n}:=\{I_{n}\leq Cn\log y_{n}\}.$
By Chebyshev’s inequality,
$\mathbb{P}(A_{n}^{c})\leq\frac{\operatorname{Var}(I_{n})}{(Cn\log y_{n}-\mathbb{E}I_{n})^{2}}\ll\frac{1}{n^{2}}.$
Let $T_{n}$ be the set of prefix configurations $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}$ satisfying $I_{n}(\boldsymbol{b})\leq Cn\log y_{n}$. Then
$\mathbb{P}\bigl((B_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}\bigr)=1-O(n^{-2}),$
and every atom in $T_{n}$ has probability at least $y_{n}^{-Cn}$. Since the total mass is at most $1$, it follows that
$|T_{n}|\leq y_{n}^{Cn}.$
Step 3: Empty-block divergence (second Borel–Cantelli). Define the empty-block event
$E_{n}:=\{B_{\ell}=0\text{ for every prime }\ell\in(y_{n},z_{n}]\}.$
Because the blocks $(y_{n},z_{n}]$ are disjoint, the events $E_{n}$ are independent. Moreover,
$\mathbb{P}(E_{n})=\prod_{\begin{subarray}{c}y_{n}<\ell\leq z_{n}\\ \ell\text{ prime}\end{subarray}}\left(1-\frac{1}{\ell-1}\right).$
Since
$1-\frac{1}{\ell-1}=\left(1-\frac{1}{\ell}\right)\left(1-\frac{1}{(\ell-1)^{2}}\right),$
Mertens’ third theorem and the convergence of $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$ imply that the product of $(1-1/\ell)$ over primes in $(y_{n},z_{n}]$ is asymptotic to $(\log y_{n})/(\log z_{n})=1/a_{n}$, since $\log z_{n}=a_{n}\log y_{n}$. Therefore
$\mathbb{P}(E_{n})\asymp\frac{\log y_{n}}{\log z_{n}}=\frac{1}{a_{n}}\asymp\frac{1}{n}.$
Therefore $\sum_{n}\mathbb{P}(E_{n})=\infty$. Since the random variables $B_{\ell}$ are mutually independent (Lemma 6.1) and the blocks $(y_{n},z_{n}]$ are pairwise disjoint, the events $E_{n}$ are mutually independent. The second Borel–Cantelli lemma therefore gives
$\mathbb{P}(E_{n}\text{ i.o.})=1.$
Step 4: Tail control. Let
$R_{n}:=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}B_{\ell}w_{\ell},\qquad F_{n}:=\{R_{n}\leq r_{n}\}.$
Since $w_{\ell}=\frac{1}{\ell}+O(\ell^{-2})$, we have
$\mathbb{E}R_{n}=\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{w_{\ell}}{\ell-1}\ll\sum_{\begin{subarray}{c}\ell>z_{n}\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell^{2}}\ll\frac{1}{z_{n}}.$
By Markov’s inequality,
$\mathbb{P}(F_{n}^{c})\leq\frac{\mathbb{E}R_{n}}{r_{n}}\ll z_{n}^{-1/2}=r_{n}.$
Thus $\sum_{n}\mathbb{P}(F_{n}^{c})<\infty$, so
$\mathbb{P}(F_{n}^{c}\text{ i.o.})=0.$
Step 5: Covering and Hausdorff measure estimate. For each $\boldsymbol{b}=(b_{\ell})_{3\leq\ell\leq y_{n}}\in T_{n}$, define
$v(\boldsymbol{b}):=\log 2+\sum_{\begin{subarray}{c}3\leq\ell\leq y_{n}\\ \ell\text{ prime}\end{subarray}}b_{\ell}w_{\ell},$
and let
$U_{n}:=\bigcup_{\boldsymbol{b}\in T_{n}}[v(\boldsymbol{b}),\,v(\boldsymbol{b})+r_{n}].$
If $A_{n}\cap E_{n}\cap F_{n}$ occurs, then the block $(y_{n},z_{n}]$ contributes nothing and the tail beyond $z_{n}$ contributes at most $r_{n}$, so $Y\in U_{n}$. Since
$\sum_{n}\mathbb{P}(A_{n}^{c})<\infty,\qquad\sum_{n}\mathbb{P}(F_{n}^{c})<\infty,\qquad\mathbb{P}(E_{n}\text{ i.o.})=1,$
the first Borel–Cantelli lemma implies that $A_{n}$ and $F_{n}$ each hold for all sufficiently large $n$ almost surely, while the second Borel–Cantelli lemma (applicable because the $E_{n}$ are independent) gives $E_{n}$ infinitely often almost surely. Hence $A_{n}\cap E_{n}\cap F_{n}$ occurs infinitely often almost surely, and on each such occurrence $Y\in U_{n}$. Therefore
$\mu_{f}\!\left(\limsup_{n\to\infty}U_{n}\right)=1.$
On the other hand,
$\sum_{n\geq 1}|T_{n}|\,r_{n}^{s}\ll\sum_{n\geq 1}y_{n}^{Cn}z_{n}^{-s/2}=\sum_{n\geq 1}y_{n}^{Cn-a_{n}s/2}\leq\sum_{n\geq 1}y_{n}^{-Cn}<\infty.$
Because $r_{n}\downarrow 0$, for every $N\geq 1$ the set $\limsup_{n\to\infty}U_{n}$ is covered by the intervals comprising $U_{n}$ for $n\geq N$, each of diameter at most $r_{N}$. Denoting by $\mathcal{H}^{s}_{\rho}(E):=\inf\bigl\{\sum_{i}(\operatorname{diam}E_{i})^{s}:E\subset\bigcup_{i}E_{i},\;\operatorname{diam}E_{i}\leq\rho\bigr\}$ the $s$-dimensional Hausdorff pre-measure at scale $\rho$ (see [14, Section 2.1]),
$\mathcal{H}^{s}_{r_{N}}\!\left(\limsup_{n\to\infty}U_{n}\right)\leq\sum_{n\geq N}|T_{n}|\,r_{n}^{s}.$
Letting $N\to\infty$ gives
$\mathcal{H}^{s}\!\left(\limsup_{n\to\infty}U_{n}\right)=0.$
Thus $\mu_{f}$ assigns full mass to a Borel set of zero $s$-dimensional Hausdorff measure, and therefore $\dim_{H}(\mu_{f})\leq s$. Since $s>0$ was arbitrary, $\dim_{H}(\mu_{f})=0$.
Finally, $\mu_{G}$ is the push-forward of $\mu_{f}$ under the map $T(u):=e^{-u}$. Since every $Z_{\ell}\geq 0$, we have $Y\geq\log 2$ almost surely, so the carrying Borel set lies in $\log 2,\infty)$. Since $T$ is Lipschitz on $\log 2,\infty)$ (with $|T^{\prime}(u)|=e^{-u}\leq\tfrac{1}{2}$), the image of any set of Hausdorff dimension 0 again has Hausdorff dimension 0 (see, e.g., [14, Corollary 2.4]). Therefore $\dim_{H}(\mu_{G})=0$. ∎
The vanishing Hausdorff dimension of the limiting measure explains why the values of $c(p)$ concentrate near arithmetic anchor points of the form $\frac{1}{2}\prod_{r\in S}(1-1/r)$, determined by the small prime divisors of p1p-1, rather than spreading uniformly over $0,\frac{1}{2}]$.
Figure 2. Empirical bin masses of $c(p)$ over the 78,49778{,}497 primes $3\leq p\leq 10^{6}$, computed via Remark 7.9 with truncation depth $J=20$, using $100$ equal-width bins covering $[0,0.5]$. The dashed bars mark selected anchor values $\varphi(m)/m$ for small even squarefree $m$. The concentration of mass near these arithmetic anchors reflects the dependence of $c(p)$ on the small prime divisors of p1p-1 and illustrates the clustering predicted by the purely singular limiting law of Corollary 5.14.
6.2. Moments and Mellin transform
Theorem 6.3 (Moment Euler product).
For every real $k>0$,
(6.1) $\mathbb{E}[X^{k}]\;=\;2^{-k}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\right).$
The product converges absolutely, with the $\ell$-th factor equal to $1+O(k/\ell^{2})$.
Proof.
Since $Y=\log 2+\sum_{\ell}Z_{\ell}$ with the $Z_{\ell}\geq 0$ independent (Lemma 6.1), the partial sums $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}\nearrow S:=\sum_{\ell}Z_{\ell}$ give $e^{-kS_{L}}\searrow e^{-kS}$. Since $0\leq e^{-kS_{L}}\leq 1$ for all $L$, by dominated convergence,
$\mathbb{E}[e^{-kS}]=\prod_{\ell\geq 3}\mathbb{E}[e^{-kZ_{\ell}}]=\prod_{\ell\geq 3}\Bigl(\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!k}\Bigr).$
Since $\mathbb{E}[X^{k}]=2^{-k}\,\mathbb{E}[e^{-kS}]$, this gives (6.1).
Writing the $\ell$-th factor as $1+a_{\ell}$ with $a_{\ell}=\bigl((\frac{\ell-1}{\ell})^{k}-1\bigr)/(\ell-1)=(e^{-kw_{\ell}}-1)/(\ell-1)$, where $w_{\ell}:=\log(\ell/(\ell-1))=1/\ell+O(1/\ell^{2})$. The bound $|e^{u}-1|\leq|u|\,e^{|u|}$ with $u=-kw_{\ell}$ gives $|a_{\ell}|\leq kw_{\ell}\,e^{kw_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, we obtain $|a_{\ell}|\leq C_{k}/\ell^{2}$ for all $\ell\geq 3$, where $C_{k}>0$ depends only on $k$. Since $\sum_{\ell}\ell^{-2}<\infty$, absolute convergence follows. ∎
For example, substituting $k=1$ and $k=2$ into (6.1) gives
$\mathbb{E}[X]=\frac{1}{2}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{2}-\ell-1}{\ell(\ell-1)},\qquad\mathbb{E}[X^{2}]=\frac{1}{4}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\frac{\ell^{3}-\ell^{2}-2\ell+1}{\ell^{2}(\ell-1)}.$
Since $X$ is supported on the compact interval $0,\tfrac{1}{2}]$ (Theorem 5.5), the Weierstrass approximation theorem implies that the moment sequence $\{\mathbb{E}[X^{k}]\}_{k=1}^{\infty}$ determines $G$ uniquely.
Theorem 6.4 (Analytic continuation).
The function
$M(s)\;:=\;2^{-s}\!\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}F_{s}(\ell),\qquad F_{s}(\ell):=\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\Bigl(\frac{\ell-1}{\ell}\Bigr)^{\!s},$
is an entire function of $s$, with the Euler product converging uniformly on compact subsets of $\mathbb{C}$. Moreover, $M(s)=\mathbb{E}[X^{s}]$ for every $s\in\mathbb{C}$, and $M(s)\neq 0$ whenever $\operatorname{Re}(s)>0$.
Proof.
We first establish the product formula $\mathbb{E}[X^{s}]=M(s)$ for every $s\in\mathbb{C}$. Write $\sigma:=\operatorname{Re}(s)$. By Lemma 6.1, $Y=\log 2+\sum_{\ell\geq 3}Z_{\ell}$ with independent $Z_{\ell}\geq 0$, and the series converges almost surely since $\sum_{\ell}\mathbb{E}[Z_{\ell}]<\infty$. Set $S_{L}:=\sum_{3\leq\ell\leq L}Z_{\ell}$ and $S:=\sum_{\ell}Z_{\ell}$. Since $Z_{\ell}\geq 0$, the partial sums are monotone: $S_{L}\leq S$ for every $L$. Hence
$|e^{-sS_{L}}|=e^{-\sigma S_{L}}\leq e^{|\sigma|S}$
for every $L$. The dominator $e^{|\sigma|S}$ is integrable: independence gives $\mathbb{E}[e^{|\sigma|S}]=\prod_{\ell\geq 3}\mathbb{E}[e^{|\sigma|Z_{\ell}}]$, and each factor equals $\frac{\ell-2}{\ell-1}+\frac{1}{\ell-1}\bigl(\frac{\ell}{\ell-1}\bigr)^{|\sigma|}=1+O_{|\sigma|}(1/\ell^{2})$, so the product converges. As $S_{L}\to S$ a.s., dominated convergence gives $\mathbb{E}[e^{-sS_{L}}]\to\mathbb{E}[e^{-sS}]$. Independence factors the finite partial products: $\mathbb{E}[e^{-sS_{L}}]=\prod_{3\leq\ell\leq L}\mathbb{E}[e^{-sZ_{\ell}}]=\prod_{3\leq\ell\leq L}F_{s}(\ell)$. Therefore $\mathbb{E}[X^{s}]=2^{-s}\,\mathbb{E}[e^{-sS}]=2^{-s}\prod_{\ell\geq 3}F_{s}(\ell)=M(s)$ for every $s\in\mathbb{C}$, extending the real-variable identity of Theorem 6.3. In particular, all negative moments $\mathbb{E}[X^{-\sigma}]$ ($\sigma>0$) are finite, a non-trivial consequence of the integrability of the dominator established above.
It remains to verify that $M(s)$ is entire and non-vanishing on $\{\operatorname{Re}(s)>0\}$. Write $w_{\ell}:=\log(\ell/(\ell-1))$ and $F_{s}(\ell)=1+a_{\ell}(s)$ with $a_{\ell}(s)=(e^{-sw_{\ell}}-1)/(\ell-1)$. The bound $|e^{z}-1|\leq|z|e^{|z|}$ gives $|a_{\ell}(s)|\leq|s|\,w_{\ell}\,e^{|s|w_{\ell}}/(\ell-1)$. Since $w_{\ell}=1/\ell+O(1/\ell^{2})$, on any compact $K\subset\mathbb{C}$ with $|s|\leq R$ we get $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$ for all large $\ell$. Since $\sum\ell^{-2}<\infty$, Weierstrass’s theorem on products of analytic functions (see e.g. [5, Ch. VII]) gives that the product converges uniformly on compact subsets of $\mathbb{C}$ to an analytic function. Each $F_{s}(\ell)$ is entire in $s$, and $2^{-s}$ is entire, so $M(s)$ is entire.
For non-vanishing with $\sigma:=\operatorname{Re}(s)>0$: each factor $F_{s}(\ell)\neq 0$ since $|F_{s}(\ell)|\geq\frac{\ell-2}{\ell-1}-\frac{e^{-\sigma w_{\ell}}}{\ell-1}$ and $e^{-\sigma w_{\ell}}<1$: at $\ell=3$, $|F_{s}(3)|\geq(1-(2/3)^{\sigma})/2>0$; at $\ell\geq 5$, $|F_{s}(\ell)|\geq(\ell-3)/(\ell-1)>0$. To conclude that the infinite product is nonzero, fix a compact $K\subset\{\operatorname{Re}(s)>0\}$. Choose $L$ so that $|a_{\ell}(s)|<1/2$ for all $\ell\geq L$ and $s\in K$ (possible since $|a_{\ell}(s)|\leq C_{R}/\ell^{2}$). For $\ell\geq L$ the principal branch $\log(1+a_{\ell}(s))$ is analytic, and $|\!\log(1+a_{\ell}(s))|\leq 2|a_{\ell}(s)|\leq 2C_{R}/\ell^{2}$, so $\sum_{\ell\geq L}\log(1+a_{\ell}(s))$ converges uniformly on $K$ and the tail product equals $\exp\!\bigl(\sum_{\ell\geq L}\log(1+a_{\ell}(s))\bigr)\neq 0$. The finitely many factors with $\ell<L$ are each nonzero, so $M(s)\neq 0$. ∎
Remark 6.5 (Zeros on the imaginary axis).
The non-vanishing assertion in Theorem 6.4 is sharp: $M(s)$ has zeros on the imaginary axis. Indeed, $F_{s}(3)=\tfrac{1}{2}+\tfrac{1}{2}(2/3)^{s}$ vanishes at $s=i\pi(2k+1)/\!\log(3/2)$ for every $k\in\mathbb{Z}$. For $\ell\geq 5$, one checks $|F_{it}(\ell)|\geq(\ell-3)/(\ell-1)>0$ for all $t\in\mathbb{R}$, so every zero of $M$ on the imaginary axis arises from the factor at $\ell=3$.
6.3. Fourier structure
Proposition 6.6 (Squared modulus of the characteristic function).
For every $\tau\in\mathbb{R}$,
$\bigl|\widehat{\mu_{f}}(\tau)\bigr|^{2}\;=\;\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\!\left(1-\frac{2(\ell-2)}{(\ell-1)^{2}}\bigl(1-\cos(\tau\log\tfrac{\ell}{\ell-1})\bigr)\right).$
Proof.
By Lemma 6.1, $|\widehat{\mu_{f}}(\tau)|^{2}=\prod_{\ell\geq 3}|\widehat{\nu_{\ell}}(\tau)|^{2}$. With $\alpha=\tau\log(\ell/(\ell-1))$ and $\widehat{\nu_{\ell}}(\tau)=\frac{\ell-2}{\ell-1}+\frac{e^{i\alpha}}{\ell-1}$,
$|\widehat{\nu_{\ell}}(\tau)|^{2}=\frac{(\ell-2)^{2}+2(\ell-2)\cos\alpha+1}{(\ell-1)^{2}}=1-\frac{2(\ell-2)}{(\ell-1)^{2}}(1-\cos\alpha),$
using $(\ell-2)^{2}+2(\ell-2)+1=(\ell-1)^{2}$. Each factor lies in $[0,1]$, so the infinite product converges in $[0,1]$; the question of whether it tends to zero as $|\tau|\to\infty$ is addressed in Section 9. ∎
Remark 6.7 (Fourier non-decay and singularity).
If $\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$, then $\mu_{f}$ is purely singular by the Riemann–Lebesgue lemma and the pure-type dichotomy (Theorem 5.9); the Lipschitz push-forward under $u\mapsto e^{-u}$ then gives singularity of $G$. Thus Fourier non-decay would provide an independent, Fourier-analytic proof of singularity.
Remark 6.8 (Numerical evidence on Fourier decay).
Numerical evaluation of $|\widehat{\mu_{f}}(\tau)|^{2}$ via the Euler product of Proposition 6.6, with the product truncated at all primes $\ell\leq 10^{7}$ (664,579664{,}579 factors), shows that the function oscillates, with local maxima near $0.14$ persisting at moderately large $\tau$. Convergence was verified by comparing partial products up to $10^{6}$, $5\times 10^{6}$, and $10^{7}$: the relative change from the second to the third cutoff is below $10^{-4}$ throughout the range tested. The arithmetic mean within each decade (over $10^{3}$ logarithmically spaced points in $1,10^{6}]$) decreases steadily: approximately $0.06$ for $\tau\in[10,10^{2}]$, declining to $0.008$ for $10^{5},10^{6}]$. This monotone decline is consistent with $|\widehat{\mu_{f}}(\tau)|^{2}\to 0$, which would mean $\mu_{f}$ is a Rajchman measure—a continuous measure whose Fourier transform vanishes at infinity (see Lyons [22]); establishing this rigorously remains open (Section 9).
Figure 3. Squared modulus $|\widehat{\mu_{f}}(\tau)|^{2}$ computed via the Euler product of Proposition 6.6 using all primes $\ell\leq 10^{7}$. The top panel uses a linear scale and the bottom panel a log-log scale. The function oscillates but its local averages decrease across successive decades of $\tau$.
6.4. Endpoint asymptotics at $1/2$
The value $1/2$ is the right endpoint of the support of $G$ and carries zero mass; in particular, $c(p)<1/2$ for every prime $p$ (Remark 5.16), even though $\varphi(p-1)/(p-1)=1/2$ for Fermat primes. The rate at which $G$ approaches its endpoint governs, in particular, the density of primes $p$ for which the reciprocal $1/c(p)$ is close to its limiting infimum $2$. We now determine this rate precisely.
Proposition 6.9 (Endpoint asymptotic at $1/2$).
Let $G$ be the limiting distribution function from Theorem 5.5, and let $X$ be a random variable with distribution $G$. Define
$\kappa:=2e^{-\gamma}\prod_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}\Bigl(1-\frac{1}{(\ell-1)^{2}}\Bigr)\approx 0.7413.$
Then, as $\varepsilon\downarrow 0$,
$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}.$
Proof.
By Lemma 6.1,
$X\overset{d}{=}\frac{1}{2}e^{-S},\qquad S:=\sum_{\begin{subarray}{c}\ell\geq 3\\ \ell\textup{ prime}\end{subarray}}Z_{\ell},$
where the $Z_{\ell}$ are independent and satisfy
$\mathbb{P}(Z_{\ell}=0)=\frac{\ell-2}{\ell-1},\qquad\mathbb{P}\!\left(Z_{\ell}=\log\frac{\ell}{\ell-1}\right)=\frac{1}{\ell-1}.$
Recall from Lemma 5.10 the notation $Q(y):=\prod_{3\leq\ell\leq y,\,\ell\textup{ prime}}(\ell-2)/(\ell-1)$. For $u>0$ one has
$\frac{u}{1+u}<\log(1+u)\leq u.$
Applying this with $u=1/(\ell-1)$ gives
$\frac{1}{\ell}<\log\frac{\ell}{\ell-1}\leq\frac{1}{\ell-1}.$
Hence, if $\ell\leq(2t)^{-1}$, then
$Z_{\ell}\neq 0\;\Longrightarrow\;Z_{\ell}>\frac{1}{\ell}\geq 2t>t,$
so the event $\{S\leq t\}$ forces $Z_{\ell}=0$ for every odd prime $\ell\leq(2t)^{-1}$. Therefore
(6.2) $\mathbb{P}(S\leq t)\leq Q\!\bigl((2t)^{-1}\bigr).$
For the lower bound, set
$z_{t}:=\frac{\log(1/t)}{t}\qquad(0<t<e^{-1}).$
By independence,
(6.3) $\mathbb{P}(S\leq t)\geq Q(z_{t})\,\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\leq t\right).$
Since $Z_{\ell}\leq 1/(\ell-1)$ almost surely,
$\mathbb{E}\!\left[\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}\right]=\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}\leq\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}\frac{1}{(\ell-1)^{2}}\ll\frac{1}{z_{t}}=\frac{t}{\log(1/t)}.$
Thus Markov’s inequality gives
$\mathbb{P}\!\left(\sum_{\begin{subarray}{c}\ell>z_{t}\\ \ell\textup{ prime}\end{subarray}}Z_{\ell}>t\right)\ll\frac{1}{\log(1/t)}=o(1),$
and (6.3) becomes
(6.4) $\mathbb{P}(S\leq t)\geq Q(z_{t})\,(1-o(1)).$
Now
$Q(y)=\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{(\ell-1)^{2}}\right).$
The second product converges to $\prod_{\ell\geq 3}(1-(\ell-1)^{-2})$, while Theorem 3.1 gives
$\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\textup{ prime}\end{subarray}}\left(1-\frac{1}{\ell}\right)\sim\frac{2e^{-\gamma}}{\log y}.$
Hence
$Q(y)\sim\frac{\kappa}{\log y}\qquad(y\to\infty).$
Since $\log((2t)^{-1})\sim\log z_{t}\sim\log(1/t)$ as $t\downarrow 0$, (6.2) and (6.4) imply
$\mathbb{P}(S\leq t)\sim\frac{\kappa}{\log(1/t)}\qquad(t\downarrow 0).$
Since $\mu_{f}$ is continuous by Proposition 5.13, the law of $S=Y-\log 2$ is also continuous, so $\mathbb{P}(S<t)=\mathbb{P}(S\leq t)$. Therefore
$1-G\!\left(\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(X>\frac{1}{2}-\varepsilon\right)=\mathbb{P}\!\left(S<-\log(1-2\varepsilon)\right).$
As $\varepsilon\downarrow 0$,
$\log(1-2\varepsilon)=2\varepsilon+O(\varepsilon^{2}),$
so $\log(1/(-\log(1-2\varepsilon)))\sim\log(1/\varepsilon)$. The stated asymptotic follows. ∎
Corollary 6.10 (Twin-prime constant).
The constant $\kappa$ equals $\mathfrak{S}_{2}\,e^{-\gamma}$, where
$\mathfrak{S}_{2}\;:=\;2\prod_{p\geq 3}\Bigl(1-\frac{1}{(p-1)^{2}}\Bigr)\;\approx\;1.3203$
is the singular series for twin primes in the sense of Hardy and Littlewood [15]. Hardy and Littlewood write $2C_{2}$ for this quantity; we use $\mathfrak{S}_{2}$ to avoid collision with the convention that reserves $C_{2}\approx 0.6602$ for $\prod_{p\geq 3}p(p-2)/(p-1)^{2}$.
Proof.
The identity $\kappa=\mathfrak{S}_{2}\,e^{-\gamma}$ follows from the factorisation $(\ell-2)/(\ell-1)=(1-1/\ell)(1-(\ell-1)^{-2})$: the first factor produces the Mertens asymptotic $2e^{-\gamma}/\!\log y$, while the product of the second converges to $\mathfrak{S}_{2}/2$. ∎
Corollary 6.11 (Endpoint modulus at $1/2$).
As $\varepsilon\downarrow 0$,
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=1-G\!\left(\frac{1}{2}-\varepsilon\right)\asymp\frac{1}{\log(1/\varepsilon)}.$
In particular, for every $\alpha>0$, one does not have
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha})\qquad(\varepsilon\downarrow 0),$
so $G$ is not Hölder of any positive order at the endpoint $1/2$.
Proof.
The two-sided order follows immediately from Proposition 6.9. If for some $\alpha>0$ one had
$G\!\left(\frac{1}{2}\right)-G\!\left(\frac{1}{2}-\varepsilon\right)=O(\varepsilon^{\alpha}),$
then Proposition 6.9 would force
$\frac{1}{\log(1/\varepsilon)}=O(\varepsilon^{\alpha}),$
which is impossible because
$\varepsilon^{\alpha}\log(1/\varepsilon)\to 0\qquad(\varepsilon\downarrow 0).$
7. Explicit Lower Bounds via the Number of Distinct Prime Factors
The explicit bounds below are deterministic and hold for every prime $p\geq 3$; their strength depends on $\omega(p-1)$. By the shifted-prime analogue of the Erdős–Kac law [12] (proved in [16, Theorem 6]), $\omega(p-1)$ is typically of order $\log\log p$. Since the $K$-th prime satisfies $p_{K}\sim K\log K$ by the prime number theorem, Mertens’ third theorem gives $B(\omega(p-1))\asymp 1/\!\log\log\log p$ for a typical prime. The sharper scale $1/\!\log\log p$ arises in Corollary 7.2 and Corollary 7.3 from the primorial comparison $\omega(p-1)\leq m(p-1)$ together with $p_{m(p-1)}\sim\log p$. Actual divisibility of p1p-1 by all primes up to $p_{m(p-1)}$ occurs only in the Linnik/primorial-progression construction used for the matching upper bound and sharpness arguments (Proposition 4.2, Theorem 7.4). For primes with unusually many prime factors in p1p-1, the bound is correspondingly weaker.
Theorem 7.1 (Explicit lower bound on $c(p)$ via the number of distinct prime factors).
Let $p\geq 3$ be a prime, and let $\omega=\omega(p-1)$ denote the number of distinct prime factors of p1p-1. Then
$c(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),\>P_{3}:=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3^{j}}\Bigr)\approx 0.5601260779$
(that is, $P_{3}=P(3)$), where $p_{1}<p_{2}<\cdots<p_{\omega}$ are the first $\omega$ primes.
Proof.
Write the distinct prime factors of p1p-1 as $r_{1}<r_{2}<\cdots<r_{\omega}$. Then
$\frac{\varphi(p-1)}{p-1}=\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{r_{i}}\Bigr)\geq\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Since $P(t):=\prod_{j=1}^{\infty}(1-t^{-j})$ is increasing in $t>1$ (each factor $1-t^{-j}$ is positive and increasing for $t>1$, and the product converges), we have $P(p)\geq P(3)=P_{3}$. Therefore, by (2.1),
$c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)\geq P_{3}\prod_{i=1}^{\omega}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
as claimed. ∎
Corollary 7.2 (Matching lower bound).
There exists an absolute constant $c>0$ such that for all sufficiently large $x$,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;\frac{c}{\log\log x}.$
Consequently,
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\asymp\;\frac{1}{\log\log x}.$
Proof.
The prime $p=2$ satisfies $c(2)=\prod_{j\geq 1}(1-2^{-j})\approx 0.2888$, a fixed positive constant that exceeds $c/\!\log\log x$ for all large $x$. For every odd prime $p\leq x$, the primorial bound $N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1<x$ gives $\omega(p-1)\leq m(x)$. By Theorem 7.1 and the fact that each factor $(1-1/p_{i})\in(0,1)$,
$c(p)\;\geq\;P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Mertens’ third theorem and Lemma 4.1 ($p_{m(x)}\sim\log x$) now yield
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\;\geq\;P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\asymp\;\frac{1}{\log\log x}.$
The matching upper bound is Proposition 4.2. ∎
Corollary 7.3 (Pointwise universal lower bound).
Recall the primorial-counting function from Section 4:
$m(n):=\max\{k\geq 1:N_{k}\leq n\}.$
Then for every prime $p\geq 3$,
$c(p)\;\geq\;P_{3}\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
In particular, there exists an absolute constant $c_{0}>0$ such that, for all sufficiently large primes $p$,
$c(p)\geq\frac{c_{0}}{\log\log p}.$
Proof.
By Theorem 7.1,
$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
Also,
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1,$
so $\omega(p-1)\leq m(p-1)$. Since each factor $1-\frac{1}{p_{i}}$ lies in $(0,1)$,
$\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr),$
which proves the first claim.
Finally, by Mertens’ third theorem and Lemma 4.1,
$\prod_{i=1}^{m(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\asymp\frac{1}{\log p_{m(p-1)}}\asymp\frac{1}{\log\log p},$
as $p\to\infty$ through the primes. This gives the second claim; equivalently, $1/c(p)\ll\log\log p$. ∎
Theorem 7.4 (Sharp constant for the extremal order).
$\limsup_{\begin{subarray}{c}p\to\infty\\ p\textup{ prime}\end{subarray}}\frac{1}{c(p)\,\log\log p}\;=\;e^{\gamma}.$
In particular, $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ for every prime $p$, and this leading coefficient is best possible.
Proof.
Upper bound. Let $p\geq 3$ be prime, with $\omega:=\omega(p-1)$ distinct prime factors $r_{1}<\cdots<r_{\omega}$ of p1p-1. Since $t\mapsto t/(t-1)$ is strictly decreasing for $t>1$ and $r_{i}\geq p_{i}$ for each $i$,
$\frac{p-1}{\varphi(p-1)}=\prod_{i=1}^{\omega}\frac{r_{i}}{r_{i}-1}\;\leq\;\prod_{i=1}^{\omega}\frac{p_{i}}{p_{i}-1}.$
The primorial bound $N_{\omega}=\prod_{i=1}^{\omega}p_{i}\leq\prod_{r\mid(p-1)}r\leq p-1$ gives $\omega\leq m(p-1)$, so extending the product (each factor exceeds $1$):
$\frac{p-1}{\varphi(p-1)}\;\leq\;\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}.$
By Mertens’ third theorem (Theorem 3.1) and Lemma 4.1,
$\prod_{i=1}^{m(p-1)}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{m(p-1)}\;\sim\;e^{\gamma}\log\log p.$
Since $1/P(p)=1+O(1/p)\to 1$, we obtain $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$.
Lower bound. For each $k\geq 1$, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, which exists by Dirichlet’s theorem. By Linnik’s theorem, $p^{*}(k)\leq C_{0}N_{k}^{\,L}$ for absolute constants $C_{0},L>0$.
Since $N_{k}\mid(p^{*}(k)-1)$, every prime $p_{i}$ with $i\leq k$ divides $p^{*}(k)-1$, and additional prime factors of $p^{*}(k)-1$ beyond $p_{k}$ only increase the product. Hence
$\frac{p^{*}(k)-1}{\varphi(p^{*}(k)-1)}\;\geq\;\prod_{i=1}^{k}\frac{p_{i}}{p_{i}-1}\;\sim\;e^{\gamma}\log p_{k}.$
For the asymptotics of $\log\log p^{*}(k)$: by Linnik’s bound, $\log p^{*}(k)\leq\log C_{0}+L\log N_{k}\sim Lp_{k}$ (using $\log N_{k}=\vartheta(p_{k})\sim p_{k}$ from Lemma 4.1), so $\log\log p^{*}(k)\leq\log p_{k}+\log L+o(1)$. Since $p^{*}(k)>N_{k}$ and $\log N_{k}\sim p_{k}$, we also have $\log\log p^{*}(k)\geq\log p_{k}+o(1)$. Therefore $\log\log p^{*}(k)\sim\log p_{k}$. Combined with $P(p^{*}(k))\to 1$:
$\frac{1}{c(p^{*}(k))\,\log\log p^{*}(k)}\;\geq\;\frac{(1+o(1))\,e^{\gamma}\log p_{k}}{(1+o(1))\,\log p_{k}}\;\to\;e^{\gamma}.$
Together, the two bounds yield $\limsup=e^{\gamma}$. ∎
Remark 7.5 (On effectivity).
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ can be made fully effective by combining explicit Mertens estimates (Rosser–Schoenfeld [29, Theorems 4 and 7], Dusart [9, Theorem 5.9]) with the Linnik-type bound $L\leq 5.18$ of Xylouris [37, Theorem 1.1], whose implied constant is effectively computable (though no numerical value has been published).
Remark 7.6 (Convergence rate of the Euler product).
Direct computation gives $P(5)\approx 0.7603$, $P(101)\approx 0.9900$, and $P(p)>1-10^{-9}$ for $p>2^{30}$; the correction factor $P(p)$ is thus negligible for all but the smallest primes.
Define $B(\omega):=P_{3}\prod_{i=1}^{\omega}(1-1/p_{i})$; by Theorem 7.1, $c(p)\geq B(\omega(p-1))$ for every prime $p\geq 3$. Table 1 lists representative values of $B(K)$ for small values of $K$.
$\omega(p-1)\leq K$ $B(K)$
11 $P_{3}/2\approx 0.280063$
55 $\approx 0.116390$
$10$ $\approx 0.0884704$
$15$ $\approx 0.0776918$
2020 $\approx 0.0715828$
Table 1. Representative values of the explicit lower bound $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ appearing in Theorem 7.1. Thus every prime $p\geq 3$ with $\omega(p-1)\leq K$ satisfies $c(p)\geq B(K)$.
Remark 7.7 (On the case $\omega(p-1)=1$).
If $\omega(p-1)=1$ then $p$ is a Fermat prime, $\varphi(p-1)/(p-1)=1/2$, and $c(p)=P(p)/2$; since $P(p)>P(3)$ for every prime $p\geq 5$, each known Fermat prime $p\geq 5$ satisfies $c(p)>B(1)$, so the bound is not tight. Along any unbounded sequence of Fermat primes, should one exist, $P(p)\to 1$ and thus $c(p)\to 1/2$.
Corollary 7.8 (Fully explicit coarse lower bound for large primes).
Let $p$ be a prime with $p>2^{30}$. Then
$c(p)>\frac{0.56}{\log p}.$
Moreover, if $\omega(p-1)\leq 10$, then
$c(p)>0.088.$
Proof.
The claim under $\omega(p-1)\leq 10$ follows from the bound $c(p)\geq B(\omega(p-1))$ (Theorem 7.1), since
$B(10)=P_{3}\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.08847>0.088.$
For the general bound, since every prime divisor of p1p-1 is at most p1p-1,
$\frac{\varphi(p-1)}{p-1}=\prod_{r\mid(p-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq p-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr).$
By Dusart’s explicit product estimate [9, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
$\prod_{\begin{subarray}{c}r\leq x\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr)\;=\;\frac{e^{-\gamma}}{\log x}\,\Bigl(1+O^{*}\!\Bigl(\frac{0.2}{\log^{3}x}\Bigr)\Bigr),$
where $O^{*}(\delta)$ denotes a quantity of absolute value at most $\delta$. Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
Also,
$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>1-10^{-9}.$
Therefore
$c(p)$ $=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)(1-10^{-9})\,\frac{e^{-\gamma}}{\log(p-1)}$
$>\;\frac{0.56}{\log p},$
using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$, and $1/(p-1)<10^{-9}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
$\frac{1}{c(p)}\;<\;\frac{\log p}{0.56}\;<\;1.79\log p.$
The $1/c(p)<1.79\log p$ bound is explicit and valid uniformly for every prime $p>2^{30}$ without any factorisation data. The sharp asymptotic scale for the extremal order is the much slower $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ of Theorem 7.4.
Remark 7.9 (Evaluation of $c(p)$).
Given the factorisation of p1p-1, the value $c(p)$ can be computed to $D$ correct decimal places (absolute error < $10^{-D}$) using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations, where $k=\omega(p-1)$ is the number of distinct prime factors of p1p-1, by computing the totient ratio $\varphi(p-1)/(p-1)$ exactly and truncating the Euler product $P(p)$ at depth $J=\lceil D\log 10/\!\log p\rceil$; writing $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, the tail satisfies $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$.
8. Cryptographic Consequences
The density $c(p)$ arises directly as a reduction constant in the PRIM-LWE variant of the learning with errors (LWE) problem [47], introduced by Sehrawat, Yeo, and Desmedt [31] to construct verifiable secret sharing schemes with “free” verification. The LWE-to-PRIM-LWE reduction proceeds by rejection-sampling: a uniformly random $n\times n$ matrix over $\mathbb{F}_{p}$ is accepted if and only if its determinant is a primitive root modulo $p$, and the acceptance probability in dimension $n$ is $c_{n}(p)$ (see Section 2). Since $c_{n}(p)$ is decreasing in $n$ with limit $c(p)$, the dimension-uniform expected rejection-sampling overhead is $1/c(p)$. This overhead represents a tightness loss in the security reduction: by the analysis of [31], the hardness guarantee for PRIM-LWE degrades by a factor of $c(p)$ relative to standard LWE, making tight bounds on $c(p)$ essential for concrete security estimates. Since $\inf_{p\textup{ prime}}c(p)=0$ (Theorem 3.3), the overhead is unbounded in general; the explicit bounds of Section 7 quantify it for any given prime.
The results of the preceding sections translate immediately into cryptographic bounds. We collect the principal consequences.
Corollary 8.1 (Compact overhead bound).
Let $P_{3}$ be as in Theorem 7.1. If $p\geq 3$ is a prime modulus with $\omega(p-1)\leq K$, then the dimension-uniform expected overhead in the LWE-to-PRIM-LWE reduction satisfies
$\frac{1}{c(p)}\;\leq\;\frac{1}{P_{3}}\prod_{i=1}^{K}\frac{p_{i}}{p_{i}-1}.$
Proof.
Immediate from Theorem 7.1 upon taking reciprocals. ∎
For concrete deployment, it is useful to evaluate $c(p)$ exactly. The following proposition upgrades the evaluation procedure described in Remark 7.9 to a certified algorithm with explicit error control.
Proposition 8.2 (Certified evaluation of $c(p)$).
Given a prime $p\geq 3$ and the complete factorisation $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
  1. (i)
    Compute the totient ratio $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(1-1/r_{i})$, which is exact as a rational number.
  2. (ii)
    Set the truncation depth $J:=\lceil D\log 10/\!\log p\rceil$ and compute the partial Euler product $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, again exact in rational arithmetic.
  3. (iii)
    Return the certified interval
    $\frac{\varphi(p-1)}{p-1}\!\left(P_{J}(p)-\frac{p^{-J}}{p-1}\right)\;\leq\;c(p)\;\leq\;\frac{\varphi(p-1)}{p-1}\,P_{J}(p).$
The width of this certified interval is at most $p^{-J}/(2(p-1))$. Moreover, in step (ii) the exact rational $P_{J}(p)$ has denominator $p^{J(J+1)/2}$, so the intermediate numerators and denominators have $O(J^{2}\log p)$ bits.
Proof.
Steps (i) and (ii) use k1k-1 and J1J-1 multiplications respectively. The certified bound in (iii) follows from the identity $P(p)=P_{J}(p)\prod_{j>J}(1-p^{-j})$ together with $\prod_{j>J}(1-p^{-j})\geq 1-\sum_{j>J}p^{-j}=1-p^{-J}/(p-1)$ and $P(p)\leq P_{J}(p)$, giving $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$. The interval width bound uses $\varphi(p-1)/(p-1)\leq 1/2$. ∎
Proposition 8.2 and the exact factorisations of p1p-1 for the NIST-standardised moduli [45, 46] yield the bounds recorded in Table 2. The factorisations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$, together with $P(p)\geq 1-1/(p-1)$, give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$. The certified evaluation yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorisation bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
Standard $p$ generic $B(\omega(p-1))$ bound sharper factorisation bound
ML-KEM (FIPS 203) $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
ML-DSA (FIPS 204) $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
Table 2. The prime moduli appearing in FIPS 203 and FIPS 204 [45, 46]. The third column gives the generic bound $B(\omega(p-1))$ from Theorem 7.1, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorisation of p1p-1 together with $P(p)\geq 1-1/(p-1)$.
Table 3 collects the overhead $1/c(p)$ for representative NTT-friendly primes from lattice-based cryptography and fully homomorphic encryption implementations.
Context $p$ $\omega(p-1)$ $1/c(p)$
ML-KEM (FIPS 203) $3329$ 22 $2.17$
ML-DSA (FIPS 204) $8380417$ 44 $3.41$
NewHope / FALCON [38, 42] $12289$ 22 $3.00$
BFV/BGV FHE (small) [40, 43, 39] $786433$ 22 $3.00$
BFV/BGV FHE (medium) [40, 43, 39] $2013265921$ 33 $3.75$
NTTRU / BLISS [44, 41] $7681$ 33 $3.75$
64-bit NTT prime [48] $2^{64}-2^{32}+1$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
Table 3. Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
Remark 8.3 (Typical overhead at cryptographic scales).
By the shifted-prime Erdős–Kac theorem (Section 7), $\omega(p-1)$ is typically of order $\log\log p$. For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes, and Table 1 gives $c(p)>0.088$. In practice, Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli. By Corollary 7.8, the overhead satisfies $1/c(p)<1.79\log p$ uniformly for every prime $p>2^{30}$ without any factorisation data; the sharp asymptotic scale is $1/c(p)\leq(e^{\gamma}+o(1))\log\log p$ (Theorem 7.4), and the order $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ is tight (Corollary 7.2).
Remark 8.4 (64-bit NTT prime).
For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ used in lattice cryptography [48], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
9. Concluding Remarks and Open Questions
This article studies the primitive-root determinant density $c(p)$ from extremal, distributional, and metric viewpoints. We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the extremal order of the reciprocal grows only on a doubly logarithmic scale, with sharp leading coefficient $e^{\gamma}$. We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a Lebesgue-null subset of $0,\tfrac{1}{2}]$ despite having full topological support. The explicit $\omega(p-1)$-based bounds make these asymptotic results effective. Section 8 translates the explicit bounds into concrete overhead estimates for the PRIM-LWE reduction in lattice-based cryptography.
We conclude with two open questions.
1. Fourier decay and the Rajchman property. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 6.6), but whether
$|\widehat{\mu_{f}}(\tau)|\to 0\qquad\text{or}\qquad\limsup_{|\tau|\to\infty}|\widehat{\mu_{f}}(\tau)|>0$
remains open. If $\widehat{\mu_{f}}(\tau)\to 0$, the measure $\mu_{f}$ would be a Rajchman measure, placing the shifted-prime distribution of $\log(n/\varphi(n))$ in the same class as many classical singular measures arising from Bernoulli convolutions. A positive $\limsup$ would instead give an independent Fourier-analytic proof of pure singularity via the Riemann–Lebesgue lemma. Numerical evidence (Remark 6.8) suggests slow decay, but the oscillatory structure—driven by the mutual incommensurability of the values $\log(\ell/(\ell-1))$ across primes $\ell$—makes rigorous conclusions difficult. Standard techniques for Bernoulli convolutions (see, e.g., Peres, Schlag, and Solomyak [26]) do not apply directly, since the contraction ratios vary with $\ell$ and do not arise from a single algebraic parameter.
2. Global modulus of continuity. Proposition 6.9 determines the behaviour of $G$ near its right endpoint: $1-G(1/2-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$, and Corollary 6.11 shows that $G$ is not Hölder of any positive order at $1/2$. Whether a logarithmic modulus of continuity holds globally—that is, whether there exists a constant $C>0$ such that $|G(\alpha)-G(\beta)|\leq C/\!\log(1/|\alpha-\beta|)$ for all $\alpha,\beta\in[0,\tfrac{1}{2}]$ with $|\alpha-\beta|$ small—remains open. For the analogous distribution of $\varphi(n)/n$ over the integers, the modulus of continuity is likewise unknown; see Schoenberg [30], Deshouillers–Hassani [7], and Toulmonde [35] for partial results in the integer case.
Data availability statement
All numerical computations are fully described by the explicit formulae in the paper and are reproducible from these descriptions alone—see Remark 5.15 for truncation and sampling parameters, and Remark 6.8 for the Fourier-decay evaluation. Python scripts reproducing Figures 13 and independently verifying Tables 13 are publicly available at https://github.com/Crypto-Vipin/primitive-root-ratio. No experimental data beyond the computations described in the text were used.